### v1

# ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM

Abstract
Multimodal Large Language Model (MLLM) often suffer from hallucinations. They over-rely on partial cues and generate incorrect responses. Recently, methods like Visual Contrastive Decoding (VCD) and Instruction Contrastive Decoding (ICD) have been proposed to mitigate hallucinations by contrasting predictions from perturbed or negatively prefixed inputs against original outputs. In this work, we uncover that methods like VCD and ICD fundamentally influence the model’s internal attention dynamics. This observation suggests that their effectiveness may not stem merely from surface-level modifications to logits but from deeper shifts in attention distribution. Inspired by this insight, we propose an attention-steerable contrastive decoding framework that directly intervenes in the model’s attention mechanisms to offer a more principled approach to mitigating hallucinations. Specifically, we introduce positive and negative steering as two complementary directions for adapting the model’s internal attention distributions. Rather than passively adjusting logits – as it is commonly done – our method dynamically modulates attention pathways within the contrastive decoding process. This enables selective enhancement or suppression of visual feature contributions in a structured manner. Furthermore, we propose a dynamic selection mechanism to identify text-centric heads – those that predominantly attend to text over visual features – for targeted positive steering, as well as a complementary mechanism to select the most critical visual tokens for negative steering, enabling more effective attention adjustments. Our experiments across multiple MLLM architectures (e.g., LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP) and diverse decoding methods (greedy search, beam search, nucleus sampling) demonstrate that our approach significantly reduces hallucinations and improves the performance on benchmarks such as POPE, CHAIR, and MMHal-Bench, while simultaneously enhancing performance on standard VQA benchmarks, including MMMU, MM-VET, ScienceQA, TextVQA, and GQA.
ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM
Yujun Wang1, Jinhe Bi2, Yunpu Ma2,3, Soeren Pirk1 1Christian-Albrechts-Universität zu Kiel (CAU) 2Ludwig Maximilian University of Munich 3Munich Center for Machine Learning
1 Introduction
Recent advances in large language models (LLMs) (Yang et al., 2024; Touvron et al., 2023; Abdin et al., 2024; Raffel et al., 2023; Brown et al., 2020; Devlin et al., 2019) have led to impressive results in a wide array of natural language processing tasks. Building on these successes, researchers have extended LLMs by visual inputs that enable multi-modal large language models (MLLMs) such as LLaVA (Liu et al., 2023b, 2024b). These MLLMs can handle complex tasks like image captioning Anderson et al. (2018), visual question answering Agrawal et al. (2016), and multimodal dialogue Das et al. (2017). Existing approaches (Dai et al., 2023; Liu et al., 2023b, 2024b; Zhou et al., 2024; Chen et al., 2023a; Alayrac et al., 2022; Bi et al., 2024) show remarkable potential to bridge the gap between vision and language.
Despite these achievements, MLLMs often inherit a critical limitation from LLMs: the tendency to produce hallucinations (Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a). These hallucinations arise when a model over-relies on partial or misleading cues, generating responses that are incorrect or do not correspond to the provided input.
Figure 1: Impact of VCD and ICD on attention distribution. We conduct an image description task using LLaVA-1.5 on 500 randomly sampled images from the COCO dataset while monitoring the internal attention distribution within the LLM component. We compare the changes in attention under different settings of Visual Contrastive Decoding (VCD), Instruction Contrastive Decoding (ICD), and their combination, relative to the original LLaVA model. The x-axis represents different attention categories: system tokens (sys), visual tokens (vis), textual tokens (text), and output tokens (output). The y-axis indicates the attention difference relative to the original model. VCD (solid blue bars) reduces attention to visual tokens while slightly increasing attention to textual tokens, with a stronger effect as the number of noising steps increases. ICD (hatched bars) exhibits a similar trend, further decreasing visual attention and increasing text attention, where stronger negative prefixes (see text in the legend) result in a more pronounced shift. When combining VCD and ICD (dotted bars), the reduction in visual attention is further amplified, while the focus on textual tokens increases. These findings indicate that the effectiveness of VCD and ICD originates from underlying shifts in the model’s attention distribution rather than solely from the contrastive decoding process.
To mitigate hallucinations, two general strategies have emerged: training-phase interventions and inference-phase interventions. In the training phase, auxiliary supervision Chen et al. (2023b) or reinforcement learning Ben-Kish et al. (2024) can help align model outputs with factual or human-preferred references. However, these approaches require additional data or complex reward modeling, which may be costly or infeasible in certain scenarios. In contrast, inference-phase methods (Zhou et al., 2024; Zhao et al., 2024; Deng et al., 2024; Wang et al., 2024a; Leng et al., 2023) aim to correct or filter erroneous outputs without retraining. Contrastive decoding is particularly appealing as it leverages negatively perturbed or prefixed inputs to steer the model away from hallucinations in a training-free manner. Two notable recent methods for contrastive decoding are Visual Contrastive Decoding (VCD) Leng et al. (2023) that perturbs an input image (e.g., via noising) to generate a “negative result” of logits, which is then subtracted from the original logits to suppress hallucinations, and second, Instruction Contrastive Decoding (ICD) Wang et al. (2024a) that prepends a negative prefix to the prompt (e.g., “You are a confused object detector”) to generate a signal that shifts the model’s predictions away from hallucinated content. Both methods offer a lightweight, yet effective approach to reducing hallucinations. However, upon closer examination, we find that these methods construct contrasting branches through surface-level modifications – either perturbing the image (VCD) or prefixing the prompt (ICD) – without explicitly addressing the underlying cause of hallucinations. Attention steering like OPERA and PAI (Liu et al., 2024c; Huang et al., 2024a) is also a common inference-phase remedy to reduce hallucination. However, PAI introduces the notion of “text inertia” – the tendency of an MLLM to keep generating text-driven content even when the image is removed – but does not articulate why steering the attention matrix is the necessary lever to overcome this inertia.
In our experiments (Fig. 1), we observe that both VCD and ICD consistently cause fundamental shifts in the internal attention distribution: they tend to reduce attention on visual tokens and amplify attention on textual tokens. This insight raises a natural question: why not directly steer the attention mechanism itself? To this end, we propose an Attention-Steerable Contrastive Decoding (ASCD) framework to manipulate attention. Specifically, the attention modification is integrated into a contrastive decoding pipeline to either enhance visual cues or to suppress negative signals. We further develop a dynamic head-selection mechanism to identify “text-centric” heads that disproportionately focus on textual cues, enabling more targeted positive adjustments. In parallel, we introduce a complementary mechanism that restricts negative steering to only the most critical visual tokens, ensuring that suppression is applied solely where necessary to mitigate hallucinations while preserving essential visual details. In summary, our contributions are as follows: (1) We analyze how recent contrastive decoding methods (VCD, ICD) create “negative samples” that fundamentally alter attention; (2) We propose an attention-steerable contrastive decoding method that explicitly modulates attention distributions to offer a more principled way to mitigate hallucinations in the inference phase; (3) We faithfully reproduce VCD and ICD to ensure fair comparison with prior work. Across three representative MLLM backbones (LLaVA-1.5 7B, LLaVA-NeXT 7B, and Phi2-SigLIP), three decoding schemes (greedy, nucleus, and beam search), and three hallucination-focused benchmarks (Rohrbach et al., 2019; Li et al., 2023b; Sun et al., 2023) (POPE, CHAIR, MMHal-Bench), our approach consistently reduces hallucinations and strengthens visual grounding. At the same time, it improves performance on standard VQA benchmarks (Yue et al., 2024; Yu et al., 2024; Lu et al., 2022; Singh et al., 2019; Hudson and Manning, 2019), including MMMU, MM-VET, ScienceQA, TextVQA, and GQA whereas other methods suffer from degraded performance on these benchmarks.
Figure 2: A motivating example of proactive attention steering in a visually ambiguous scenario. (a) shows the conversation context where the “orange” is actually tinted blue. (b) shows how the logits vary based on negative-steering. (c) shows how the logits vary based on positive-steering. (d) illustrates how attention-steerable contrastive decoding, which combines both negative and positive steering in a unified framework, reduce hallucinations and produce perception-driven answers.
2 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) have significantly advanced the field of artificial intelligence by integrating vision and language understanding, enabling a wide range of vision-language tasks (Dai et al., 2023; Zhu et al., 2023; Liu et al., 2024b, 2023b; Alayrac et al., 2022; Chen et al., 2023a; Zhou et al., 2024; Zhang et al., 2023a; Rong et al., 2025; Chen et al., 2025a; Yu et al., 2025; Guan et al., 2025; Huang et al., 2024c, 2025; Liu et al., 2025; Zhao et al., 2025b). These models typically follow a two-stage training paradigm: (1) large-scale pretraining on web-scale image-text pairs (Liu et al., 2023b; Li et al., 2023a) to learn cross-modal representations, and (2) visual instruction tuning (Liu et al., 2023a; Bi et al., 2025) on task-specific datasets to enhance multimodal instruction-following capabilities. While this paradigm has led to substantial improvements in vision-language reasoning, MLLMs still face key challenges, such as hallucination – where the model generates content that is inconsistent with the given visual input. (Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a).
Mitigating Hallucinations in MLLMs. Hallucination in MLLMs is particularly pronounced in open-ended generation tasks, where models may produce content that is not aligned with the provided visual input (Huang et al., 2024a; Jing et al., 2024; Zhang et al., 2023b). Some approaches focus on the mitigation of data bias, scaling-up of vision resolution, and alignment optimization. Lovenia et al. (2024) introduce a technique that mines 95,000 negative samples by replacing original categories, attributes, or quantity information with similar but incorrect alternatives. This fine-grained approach effectively enriches the contrastive signal during training, thereby enhancing the model’s robustness. Chen et al. (2024) propose InternVL, which scales the vision encoder up to 6 billion parameters and processes images with widths ranging from 1,664 to 6,144 pixels. While this method improves visual detail and alignment, it requires significant computational resources for pretraining with large-scale data. Sun et al. (2023) employ Reinforcement Learning from Human Feedback (RLHF) Stiennon et al. (2022) to align different modalities during training. This optimization strategy leads to a reduction in hallucinations by better integrating visual and textual cues. Bi et al. (2024) propose a representation steering method that effectively mitigates hallucination in multimodal models.
Contrastive Decoding Approaches. Recent work has explored contrastive decoding as an effective, training-free means to mitigate hallucinations Xiao et al. (2025). For instance, Leng et al. (2023) introduced Visual Contrastive Decoding (VCD), which perturbs the input image to generate a negative logit branch that is subtracted from the original predictions, while Wang et al. (2024a) employs a negative prompt to steer outputs away from hallucinated content. Huo et al. (2024) leverages a Context and Text-aware Token Selection (CT2S) strategy to selectively retain the most informative vision tokens in early decoder layers, thereby amplifying beneficial multimodal context and suppressing spurious hallucinations.
3 Preliminaries
Modern MLLMs integrate text and visual inputs based on powerful encoders that enable merging the modalities into a unified representation that is processed by a multi-layer Transformer. While these models enable producing coherent responses, they heavily rely on internal attention mechanisms that dictate how visual and textual cues are combined. As discussed in Section 3.2, subtle variations in these attention distributions can significantly impact the generated output. This observation motivates our approach: by explicitly modulating attention, we aim to enhance visual grounding and mitigate hallucinations.
3.1 MLLM Formulation
We consider a multimodal large language model (MLLM) that processes an image $\mathbf{I}$ and a text prompt $\mathbf{x}=\{x_{1},\dots,x_{N}\}$ to generate an output sequence $\mathbf{y}=\{y_{1},\dots,y_{M}\}$ in an autoregressive manner. Let $\theta$ denote the model parameters. Formally, the model maximizes:
$\mathbf{y}^{*}=\arg\max_{\mathbf{y}}\prod_{t=1}^{M}p_{\theta}\Bigl{(}y_{t}\,\Big{|}\,\mathbf{I},\mathbf{x},y_{<t}\Bigr{)},$ (1)
where $y_{<t}$ denotes all previously generated tokens.
Embeddings. A unified input is obtained from encoded image and embedded text:
$\mathbf{Z}=[f_{v}(\mathbf{I});f_{t}(\mathbf{x})].$ (2)
Transformer Architecture. The MLLM processes $\mathbf{Z}$ through $L$ Transformer layers Vaswani et al. (2023):
$\mathbf{H}^{(l)}=\mathrm{TransformerLayer}^{(l)}\!\bigl{(}\mathbf{H}^{(l-1)}\bigr{)},\quad\mathbf{H}^{(0)}=\mathbf{Z}.$ (3)
Output Prediction. The final hidden state $\mathbf{h}_{t}^{(L)}$ is mapped to a probability distribution over the vocabulary:
$p_{\theta}(y_{t}\mid\mathbf{I},\mathbf{x},y_{<t})=\mathrm{Softmax}\!\Bigl{(}\mathbf{h}_{t}^{(L)}W^{P}\Bigr{)},$ (4)
where $W^{P}$ is the output projection matrix.
3.2 Proactive Steering of Attention
In Figure 1, we show how Visual Contrastive Decoding (VCD) and Instruction Contrastive Decoding (ICD) indirectly alter attention distributions. Building on this insight, we now ask: what if we explicitly steer the model’s attention? Figure 2 provides a motivating example, illustrating how actively modulating attention can influence the final logits distribution.
Consider a simple query: “What is the color of the orange here?” The conversation context (Figure 2a) is based on LLaVA-1.5 7B, with a provided image in which the “orange” fruit appears to be tinted blue. We experiment with two distinct attention-steering scenarios: negative-steered logits (Figure 2b) and positive-steered logits (Figure 2c). In each case, we proportionally adjust the visual or textual attention before finalizing the output distribution.
In the negative-steered branch, we reduce attention to visual tokens or boost attention to the textual tokens. As shown in the histogram of logits, the model reduces its reliance on the visual input, causing it to fall back more heavily on the LLM’s inherent priors. As a result, it is more likely to generate answers that align with typical linguistic associations rather than the actual content of the image – insisting the color is “orange”. Conversely, the positive-steered branch increases attention to visual tokens or downgrades textual tokens, making the model more sensitive to the actual (albeit unexpected) color in the image. This leads the model to answer “blue” with higher probability.
In addition to these unidirectional adjustments, we further integrate attention steering into the contrastive decoding framework. Instead of using the original logits directly (as in VCD or ICD), we inject the attention-modulated logits. Mathematically, we redefine the contrastive decoding formulation by replacing the original logits adjustment with a positively steered version:
$p_{\theta}^{\text{final}}\;=\;(1+\alpha)p_{\theta}^{\text{pos-steered}}-\alpha p_{\theta}^{\text{neg-steered}},$ (5)
where $p_{\theta}^{\text{pos-steered}}$ and $p_{\theta}^{\text{neg-steered}}$ represent the output logits modified by positively or negatively steered attention.
By integrating contrastive decoding with explicit attention manipulation, our attention-steerable contrastive decoding framework (Figure 2d), sharpens the output distribution which enhances the likelihood of the correct response, while reducing the impact of competing distractors.
Figure 3: Distribution of text-centric heads across different models and experiment settings. Each heatmap visualizes how frequently a given head occurs among the most text-focused heads. The panel in the center (a) show the result of LLaVA-1.5 with a generation length of 64 tokens; (b) and (c) show results of the same model with longer generation (512 tokens) and a different image set. Despite these changes, LLaVA-1.5 exhibits minimal JS divergence, which indicates consistent text-centric heads. In contrast, Phi2-SigLIP (d) and LLaVA-NeXT (e) deviate substantially from LLaVA-1.5, revealing model-specific attention biases and higher JS divergence.
3.3 Text-centric Heads
Previously, we have highlighted the impact of adjusting attention. In this section, we discuss which heads in the model are most prone to over-reliance on textual cues. To this end, we conduct an experiment to identify "text-centric" heads, i.e., those with disproportionately high text-to-visual attention ratios, and examine their consistency under different generation conditions and image sets. The experiment setup is detailed in Appendix A.
Results and Observations. Figure 3 shows the resulting heatmaps $F$ for multiple models and generation settings. The panel in the center (a) corresponds to LLaVA-1.5 on $N=500$ images with a generation length of 64 tokens. The two heatmaps at the bottom show results of the same model but with either an increased generation length to 512 tokens (b, bottom left), or using a different set of 500 images (c, bottom right). Despite these changes, the distribution of top text-focus heads remains visually similar, and the small Jensen–Shannon (JS) divergences confirm that these text-centric heads are largely invariant under different sampling conditions for the same model.
In contrast, the Phi2-SigLIP (d, top-left) and LLaVA-NeXT (e, top-right) panels deviate significantly from LLaVA-1.5 even under the same experiment settings, with higher JS divergence. This suggests that each model has its own unique set of heads that consistently favor textual attention over visual cues. However, within a single model, the text-centric heads persist across varied prompts, image sets, and generation lengths.
Implications. The consistent presence of the text-centric heads within the same model indicates that certain heads are inherently prone to focusing on textual signals rather than visual content. In Section 4.2 we describe how this insight can be leveraged to selectively target the problematic heads when applying our positive steering strategy. Rather than uniformly amplifying attention across all heads, we concentrate on those that are most responsible for text-dominant attention, thereby avoiding unnecessary modifications to heads that are well-balanced in their visual-textual focus.
4 Methodology
Figure 4: Illustration of positive and negative steering. Left: text-centric heads are boosted (positive_steer) to emphasize visual content; Right: a small set of critical visual tokens is suppressed (negative_steer), inducing a stronger contrastive effect. These selective adjustments work in tandem to reduce hallucinations and improve grounding.
In this section, we present our attention-steerable contrastive decoding framework, which explicitly modulates the model’s attention to mitigate hallucinations. Our approach has two stages: (1) Text-centric Head Selection, which identifies the heads most prone to text-centric bias, and (2) Attention Steering, where we apply positive steering to text-centric heads and negative steering to a small subset of visually critical tokens. We then integrate these adjusted logits for generation into a contrastive decoding pipeline.
4.1 Text-centric Head Selection
Algorithm 1 Text-centric Head Selection (Offline)
1:Reference dataset $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$, MLLM with $L$ layers and $H$ heads per layer, Final text-centric head count $\kappa_{\textsc{tch}}$
2:$\mathcal{H}_{\textsc{pos}}$ (set of selected text-centric heads)
3:Initialize a global counter $F\in\mathbb{R}^{L\times H}$ to zeros
4:for $i\leftarrow 1$ to $N$ do
5:     Run MLLM on image $\mathbf{I}_{i}$ (e.g., image description)
6:     for all head $(r,c)$ in layer-head grid do
7:         Compute attention ratio:
$Q_{i}(r,c)=\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)}$
8:     end for
9:     Identify top-32 indices of $Q_{i}$ (largest ratios) and store in $\mathcal{I}_{i}$
10:     for all $(r,c)\in\mathcal{I}_{i}$ do
11:         $F(r,c)\leftarrow F(r,c)+1$
12:     end for
13:end for
14:Sort all heads $(r,c)$ in descending order by $F(r,c)$
15:Select top $\kappa_{\textsc{tch}}$ heads:
$\mathcal{H}_{\textsc{pos}}\leftarrow\text{first }\kappa_{\textsc{tch}}\text{ heads in sorted list}$
16:return $\mathcal{H}_{\textsc{pos}}$
As detailed in Algorithm 1, we start by identifying the most text-centric heads using a small reference dataset (e.g., 500 images) for a task (e.g., image description). For each sample, we compute the ratio of textual attention to visual attention (Eq. 6 in Appendix A) and take the top 32 heads with the highest ratio. We accumulate these counts over all samples, then choose the top $\kappa_{\textsc{tch}}$ heads as “text-centric”. This step is motivated by our finding (Section 3.3) that certain heads consistently favor textual content over visual cues.
Algorithm 2 Attention-Steerable Contrastive Decoding
1:Image $\mathbf{I}$, Text-centric heads $\mathcal{H}_{\textsc{pos}}$ (from Algorithm 1), Critical vis-token count $\kappa_{\textsc{vis}}$, Steer strengths $\alpha_{\textsc{pos}},\alpha_{\textsc{neg}}$, Contrastive weight $\alpha$, Truncation threshold $\beta$, MLLM with $L$ layers and $H$ heads per layer
2:$p_{\theta}^{\text{final}}$ (final logits from ASCD)
3:Step 1: Forward Pass with Positive Steering
4:for $l\leftarrow 1$ to $L$ do
5:     for $h\leftarrow 1$ to $H$ do
6:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
7:         if $(l,h)\in\mathcal{H}_{\textsc{pos}}$ then
8:              $\mathbf{A}^{(l)}_{h}\;\leftarrow\;\mathbf{A}^{(l)}_{h}\;+\;\alpha_{\textsc{pos}}\,|\mathbf{A}^{(l)}_{h}|$
9:         end if
10:     end for
11:     Normalize $\mathbf{A}^{(l)}$ and continue
12:end for
13:Obtain logits $p_{\theta}^{\text{pos-steered}}$
14:Step 2: Forward Pass with Negative Steering
15:for $l\leftarrow 1$ to $L$ do
16:     for $h\leftarrow 1$ to $H$ do
17:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
18:         Identify top-$\kappa_{\textsc{vis}}$ critical visual tokens $\mathcal{V}$
19:         for all $v\in\mathcal{V}$ do
20:              $\mathbf{A}^{(l)}_{h}(v)\;\leftarrow\;\mathbf{A}^{(l)}_{h}(v)\;-\;\alpha_{\textsc{neg}}\,|\mathbf{A}^{(l)}_{h}(v)|$
21:         end for
22:     end for
23:     Normalize $\mathbf{A}^{(l)}$ and continue
24:end for
25:Obtain logits $p_{\theta}^{\text{neg-steered}}$
26:Step 3: Contrastive Decoding with Truncation
27:$p_{\theta}^{\text{raw}}\;\leftarrow\;(1+\alpha)\,p_{\theta}^{\text{pos-steered}}\;-\;\alpha\,p_{\theta}^{\text{neg-steered}}$
[…]
29:$p_{\theta}^{\text{final}}\;\leftarrow\;p_{\theta}^{\text{raw}}.\text{masked\_fill}(p_{\theta}^{\text{pos-steered}}<\text{cutoff},-\infty)$
30:return $p_{\theta}^{\text{final}}$
4.2 Attention Steering
Text-centric Head Awareness and Critical Visual Token Selection. As shown in Figure 4, we refine our method by incorporating text-centric head selection for positive steering and critical token identification for negative steering. Specifically, given the selected text-centric heads, we positively steer them by increasing their attention weights with a strength of $\alpha_{\textsc{pos}}$. Figure 5a highlights how targeted steering in text-centric heads improves the positive steering effectiveness. Simultaneously, we apply negative steering to the top $\kappa_{\textsc{vis}}$ most critical visual tokens – those receiving the highest aggregate attention across heads – reducing their attention scores by $\alpha_{\textsc{neg}}$ across all heads. Through this strategy, we deliberately obscure only the most pivotal cues – this targeted suppression is sufficient to induce a strong hallucination effect in the negative branch, leading to improved contrastive decoding compared to a blanket suppression of all visual tokens. In Figure 5b, we demonstrate the impact of selectively applying negative steering to critical visual tokens.
Figure 5: Comparative effectiveness of selective attention steering. (a): Positive steering applied only to text-centric heads (32 heads with the highest text-to-visual ratio) outperforms random or blanket head selection across various decoding strategies (Greedy, Nucleus, Beam), leading to higher POPE Accuracy and F1. (b): Negative steering focused on a small subset of critical visual tokens, integrated with contrastive decoding, significantly reduces CHAIR metrics (less hallucination) and boosts POPE metrics compared to randomly suppressing visual tokens of the same number. These results validate that targeted attention modulation on text-centric heads (for positive steering) and critical visual tokens (for negative steering) yields stronger hallucination mitigation and more grounded responses.
Integration with Contrastive Decoding with Truncation. Building on the attention-steering process, we first obtain two output distributions: $p_{\theta}^{\text{pos-steered}}$ from the positively steered branch and $p_{\theta}^{\text{neg-steered}}$ from the negatively steered branch. We then combine these into contrastive decoding with a truncation mechanism, as detailed in the Step 3 of Algorithm 2. This process not only reinforces visually grounded predictions but also effectively mitigates the influence of spurious textual biases.
5 Experiments
To evaluate the effectiveness of our attention-steerable contrastive decoding framework in mitigating hallucinations in MLLMs, we conduct a range of experiments. This includes three diverse benchmarks – CHAIR, POPE, and MMHal-Bench – each designed to assess different aspects of object hallucinations. To ensure the broad applicability and robustness of our approach, we also test it on three representative models – LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, and employ three different decoding strategies: greedy search, nucleus sampling, and beam search. Details of the experimental settings are provided in Appendix B. Furthermore, we evaluate performance on standard VQA benchmarks including MMMU, MM-VET, ScienceQA, TextVQA, and GQA to verify that the proposed method preserves – rather than diminishes – the model’s original visual understanding.
Figure 6: Radar charts of MMHal-Bench results. Each axis represents a different evaluation dimension in MMHal-Bench, and a larger enclosed area indicates better overall performance.
Model Decoding Method CHAIRs (↓) CHAIRi (↓)
LLaVA-1.5 7B greedy Orig 53.2 13.5
VCD 56.8 15.2
ICD 52.8 13.2
ASCD 35.6 8.6
nucleus Orig 59.0 17.4
VCD 59.8 16.6
ICD 57.4 15.6
ASCD 43.6 11.3
beam Orig 54.8 15.3
VCD 58.8 16.4
ICD 52.6 13.9
ASCD 40.8 10.1
LLaVA-NeXT 7B greedy Orig 31.6 7.5
VCD 37.2 9.7
ICD 32.8 8.4
ASCD 21.8 7.0
nucleus Orig 30.4 8.0
VCD 40.4 10.4
ICD 39.4 9.9
ASCD 21.2 6.7
beam Orig 34.0 8.5
VCD 36.6 9.1
ICD 31.8 7.6
ASCD 21.0 6.5
Phi2-SigLIP greedy Orig 29.0 6.9
VCD 39.4 9.6
ICD 33.4 7.7
ASCD 21.8 5.4
nucleus Orig 36.0 9.8
VCD 36.0 8.1
ICD 37.0 9.4
ASCD 26.0 8.0
beam Orig 30.4 6.9
VCD 36.0 8.4
ICD 31.0 7.0
ASCD 24.6 5.7
Table 1: CHAIR Evaluation Results. Lower CHAIRs and CHAIRi values indicate better performance in reducing hallucination. The best values for each metric within a model-decoding combination are highlighted in bold.
It is important to note that current benchmarks for evaluating multimodal models are highly variable. For example, baseline models such as LLaVA-1.5 7B often report different metric values between different papers. Moreover, the CHAIR metric relies on random image sampling, which further complicates direct comparisons between papers. To address these issues, we faithfully reproduced both VCD and ICD using the parameters specified in their original papers and repositories, ensuring that our evaluations are conducted under consistent conditions. This allows for a more reliable comparison between our method and existing approaches.
Model Decoding Method Popular Acc Popular F1 Random Acc Random F1 Adv. Acc Adv. F1 Avg. Acc Avg. F1
LLaVA-1.5 7B greedy Orig 85.83 84.35 86.67 85.51 83.60 82.31 85.37 84.06
VCD 84.67 83.59 86.43 85.53 81.70 80.93 84.27 83.35
ICD 83.47 80.85 83.68 81.51 82.07 79.56 83.07 80.64
PAI - - - - - - 85.82 85.79
ASCD 87.20 86.69 89.48 89.09 82.90 82.97 86.53 86.25
nucleus Orig 83.47 81.74 84.78 83.44 80.83 79.53 83.03 81.57
VCD 83.83 82.61 85.12 84.12 80.97 80.18 83.31 82.30
ICD 82.63 79.85 82.68 80.43 81.07 78.58 82.13 79.62
PAI - - - - - - 81.72 82.87
ASCD 86.47 85.56 87.94 87.26 82.83 82.39 85.75 85.07
beam Orig 85.87 84.39 86.70 85.55 83.63 82.36 85.40 84.10
VCD 84.43 83.30 86.19 85.31 82.20 81.30 84.27 83.30
ICD 83.47 80.83 83.61 81.43 82.03 79.51 83.04 80.59
PAI - - - - - - 86.33 85.89
ASCD 87.20 86.69 89.48 89.09 82.87 82.95 86.52 86.24
LLaVA-NeXT 7B greedy Orig 83.97 81.77 85.09 83.26 82.73 80.64 83.93 81.89
VCD 84.87 83.12 86.19 84.84 83.53 81.89 84.86 83.28
ICD 84.53 82.63 85.70 84.15 83.10 81.33 84.44 82.70
ASCD 84.90 83.30 86.39 85.09 83.27 81.82 84.85 83.40
nucleus Orig 81.73 79.26 83.61 81.75 79.87 77.81 81.74 79.61
VCD 84.20 82.51 84.78 83.26 81.67 80.07 83.55 81.95
ICD 83.60 81.68 85.29 83.78 82.13 80.47 83.67 81.98
ASCD 84.60 82.86 86.19 84.77 83.27 81.65 84.69 83.09
beam Orig 84.17 82.04 85.26 83.49 82.90 80.88 84.11 82.14
VCD 84.67 82.86 86.19 84.81 83.13 81.42 84.66 83.03
ICD 84.57 82.68 85.74 84.19 83.13 81.37 84.48 82.75
ASCD 84.97 83.39 86.43 85.14 83.33 81.91 84.91 83.48
Phi2-SigLIP greedy Orig 87.10 85.95 88.45 87.57 86.03 84.97 87.19 86.16
VCD 86.00 85.14 87.97 87.37 84.70 84.09 86.22 85.53
ICD 85.50 84.14 87.25 86.15 84.73 83.44 85.83 84.58
ASCD 87.77 86.74 88.90 88.14 86.77 85.81 87.81 86.90
nucleus Orig 85.73 84.49 86.87 85.87 83.93 82.96 85.51 84.44
VCD 85.60 84.72 86.91 86.21 84.30 83.64 85.60 84.86
ICD 84.90 83.46 85.98 84.86 83.00 81.73 84.63 83.35
ASCD 87.50 86.41 88.52 87.69 86.33 85.29 87.45 86.46
beam Orig 87.10 85.95 88.45 87.57 86.03 84.97 87.19 86.16
VCD 86.43 85.63 87.90 87.36 84.57 83.92 86.30 85.64
ICD 85.50 84.14 87.25 86.15 84.73 83.44 85.83 84.58
ASCD 87.77 86.74 88.90 88.14 86.77 85.81 87.81 86.90
Table 2: POPE Evaluation Results. The best values for each metric within a model-decoding combination are highlighted in bold. If our ASCD achieves the second-best result, it is additionally marked with an underline.
CHAIR. Table 1 shows the CHAIR metrics (CHAIRs and CHAIRi), which measure object hallucination in image captioning. Across all models and decoding strategies, ASCD consistently achieves lower CHAIR values than Orig, VCD, or ICD, which illustrates ASCD’s effectiveness at mitigating object-level hallucinations.
POPE. Table 2 reports the accuracy and F1 scores under the POPE evaluation, which probes object presence with random, popular, and adversarial queries. Higher values indicate fewer hallucinations. Again, ASCD achieves the best or near-best performance in all cases. These gains persist across different model architectures, suggesting that attention steering is robust to model size and design variations.
MMHal-Bench. Figure 7 illustrates the radar charts of MMHal-Bench results for LLaVA-1.5 7B under greedy and nucleus decoding. Each axis represents a sub-dimension of the benchmark, and a larger area signifies better overall performance. ASCD exhibits the largest enclosed area, outperforming baseline, VCD, and ICD in most dimensions. This improvement aligns with the CHAIR and POPE findings, underscoring the benefit of selectively steering attention to reduce hallucinations.
Standard VQA Benchmarks. To verify that ASCD does not sacrifice a model’s general visual-question-answering ability, it’s evaluated on five widely-used VQA datasets. Across all three backbones and all decoding strategies, ASCD either matches or surpasses the original model on every dataset, while VCD and ICD consistently degrade performance as shown in the Appendix C.
Summary. Our experiments confirm that ASCD effectively reduces hallucinations and improves alignment with visual content, regardless of the model or decoding strategy employed.
6 Conclusion
We have shown that existing contrastive methods (e.g. VCD and ICD) inadvertently shift the internal attention distribution in MLLM, prompting us to investigate a more direct and principled way to modulate attention. We proposed an attention-steerable contrastive decoding framework that positively steers text-centric heads while negatively steering only the most critical visual tokens.
Our method consistently reduces hallucinations on CHAIR, POPE, and MMHal-Bench, outperforming both baseline and previous contrastive approaches with improved and uncompromised general VQA capability. By targeting precisely those heads and tokens, we effectively mitigate spurious textual biases while preserving essential visual context.
Limitations
A key limitation of ASCD is its incompatibility with FlashAttention. Because our method must dynamically modify the attention matrix at inference time, it cannot make use of the fused kernels, leading to higher memory consumption and slower decoding. A promising workaround is to distill the steering signal into the model during training: we can add an auxiliary loss – e.g., a KL-divergence term – that drives the native attention distribution to approximate the ASCD target distribution. If successful, the model would internalise the hallucination-mitigation behaviour, removing the need for on-the-fly edits and restoring FlashAttention speed-ups. We regard training-time attention regularization as a promising direction: it could distill the hallucination-robust behaviour discovered by training-free, attention-modified methods into the model itself, so that at inference the model retains this robustness while fully benefiting from FlashAttention’s speed and memory efficiency.
Appendix A Text-centric Heads Experiment Settings
The following setup applies to Section 3.3.
We select $N=500$ images $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$ (from COCO) and run an MLLM (LLaVA-1.5) in an image description task. During each generation, we track the ratio of textual attention to visual attention for every head:
$Q_{i}\in\mathbb{R}^{R\times C},\quad Q_{i}(r,c)\;=\;\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)},$ (6)
where $r$ and $c$ index each head (for instance, $R=C=32$). We then identify the top-$k$ heads with the highest ratio values and mark them in a binary mask:
$M_{i}(r,c)\;=\;\begin{cases}1,&\text{if }(r,c)\in\text{top-$k$ indices of }Q_{i},\\ 0,&\text{otherwise.}\end{cases}$ (7)
Finally, we aggregate these masks across all $N$ images:
$F\;=\;\sum_{i=1}^{N}M_{i},$ (8)
so that $F(r,c)$ records how frequently head $(r,c)$ appears among the most text-focused heads.
Appendix B Evaluation Settings
B.1 Baseline Models and Decoding Methods
We evaluate our proposed approach using three representative models and multiple decoding strategies to demonstrate its broad applicability and robustness.
LLaVA-1.5 7B is a minimalist yet efficient model that has served as the foundation for extensive studies in large multimodal modeling. LLaVA-NeXT 7B builds on LLaVA-1.5 with improvements in visual reasoning, higher input resolution, and enhanced world knowledge, resulting in superior performance on several benchmarks. Phi2-SigLIP leverages the Phi-2 backbone and a SigLIP-based vision tower, and is trained on the ShareGPT4V dataset, offering a compact alternative with competitive capabilities.
To assess the reliability and generalizability, we experiment with three decoding strategies: greedy search, nucleus sampling, and beam search. Greedy decoding yields deterministic outputs, while nucleus sampling and beam search enable for more diverse generation.
B.2 Datasets
With the widespread adoption of deep learning across computer vision and natural language processing Lu et al. (2023); Chen et al. (2025b, a); Cui et al. (2025); Yuan et al. (2025); Lu et al. (2024a, b); Zhao et al. (2025a, c); Xu et al. (2024, 2025), large language models (LLMs) have rapidly become foundational tools in AI systems Wang et al. (2024b); Ma et al. (2025); Wang et al. (2025b, a); Zeng et al. (2024); Xin et al. (2025); Du et al. (2021, 2025); Zhang et al. (2023c). However, despite their remarkable capabilities, MLLMs remain susceptible to hallucination—generating descriptions inconsistent with the input image. This phenomenon has sparked growing interest in evaluation benchmarks that specifically target hallucination in image-grounded generation.
We evaluate our approach on three hallucination-targeted benchmark datasets designed to probe object hallucination in multimodal large language models.
CHAIR. The Caption Hallucination Assessment with Image Relevance (CHAIR) metric quantifies the degree of hallucination in generated captions by measuring the fraction of objects mentioned that do not actually appear in the image. It is computed at both the instance-level (CHAIRi) and the sentence-level (CHAIRs), offering insight into how well a caption adheres to veridical image content.
POPE. The Polling-based Object Probing Evaluation (POPE) assesses hallucination by querying the model with binary questions (e.g., “Is there a car in the image?”). By balancing queries about present and absent objects, and using different sampling strategies (random, popular, adversarial), POPE effectively reveals the influence of language priors on model predictions. This method provides a robust measure of object hallucination across multiple datasets such as MSCOCO, A-OKVQA, and GQA.
MMHal-Bench. MMHal-Bench is a new evaluation benchmark specifically designed to challenge large multimodal models in hallucination. Comprising 96 difficult questions based on images, along with detailed ground-truth answers and image content annotations, MMHal-Bench offers a comprehensive testbed to assess the model’s ability to produce veridical responses in ambiguous visual scenarios.
We also evaluate our approach on five standard VQA benchmarks to verify that our method does not compromise, and may even enhance, the original VQA capabilities.
MM-VET. The MM-VET benchmark evaluates large multimodal models on integrated vision–language abilities that go beyond conventional VQA. It defines six core capabilities—recognition, OCR, knowledge, language generation, spatial awareness, and math—and constructs 16 task types that require their various combinations. MM-Vet comprises 200 images paired with 218 open-ended questions; answers are graded by a GPT-4-based evaluator that yields unified scores, enabling fine-grained, per-capability diagnostics.
ScienceQA. The ScienceQA benchmark targets multimodal, curriculum-level science reasoning. It contains multiple-choice questions drawn from elementary- to high-school materials, each optionally paired with text passages, natural or diagrammatic images, or both. In this study, answers are graded for accuracy.
GQA. The GQA benchmark targets real-world compositional visual reasoning. Questions span objects, attributes, relations, spatial logic, comparisons, and multi-hop inference, making GQA a stringent yardstick for models that aspire to deep vision-language understanding rather than pattern-matching.
TextVQA. The TextVQA benchmark focuses on visual question answering that explicitly requires reading and reasoning over scene text. It contains 45,336 open-ended questions paired with 28,408 everyday images drawn from the Open Images dataset. Questions span practical tasks such as reading product labels, traffic signs, receipts, and clocks, so models must combine OCR, spatial grounding, and commonsense reasoning.
MMMU. The Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark gauges college-level, expert-style reasoning across six broad disciplines—Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering.
B.3 Hyperparameters
The experiments are conducted with the following hyperparameter settings to obtain the best result:
$\kappa_{\textsc{vis}}=0.1$, $\alpha_{\textsc{neg}}=1.0$
LLaVA-1.5: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.6$, $\alpha=1.0$, and $\beta=0.5$.
LLaVA-NeXT: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.7$, $\alpha=1.0$, and $\beta=0.7$.
Phi2-SigLIP: $\kappa_{\textsc{tch}}=128$, $\alpha_{\textsc{pos}}=0.8$, $\alpha=0.5$, and $\beta=0.1$.
Appendix C Standard VQA Capability
Our approach substantially curbs hallucinations through explicit attention steering while simultaneously enhancing performance on standard VQA tasks, a trade-off that competing methods typically fail to avoid.
C.1 MM-VET
Method MM-VET Score
Orig 31.2
VCD 30.3
ICD 33.2
ASCD 33.2
Table 3: MM-VET scores for different methods.
C.2 ScienceQA
Method SQA Score
Orig 67.55
VCD 67.55
ICD 67.32
ASCD 69.51
Table 4: SQA scores for different methods.
C.3 GQA
Method GQA Score
Orig 61.28
VCD 59.38
ICD 59.99
ASCD 61.27
Table 5: GQA scores for different methods.
C.4 TextVQA
Method TextVQA Score
Orig 57.82
VCD 55.07
ICD 57.66
ASCD 57.91
Table 6: TextVQA scores for different methods.
C.5 MMMU
Method MMMU Score
Orig 0.342
VCD 0.333
ICD 0.360
ASCD 0.348
Table 7: MMMU scores for different methods.
Appendix D Visualization Example
Figure 7: Oringinal LLaVA-1.5 drifts completely, describing a riverside sidewalk, parked cars, and a stroller that do not exist in the image. VCD and ICD correctly identify the beach and ATV but still hallucinate extra objects such as a surfboard and a boat (highlighted in yellow). Our ASCD caption mentions only entities that are actually present and omits non-existent items, illustrating the improved visual grounding achieved through attention-steerable contrastive decoding.

### v2

@@
 Abstract
-Multimodal Large Language Model (MLLM) often suffer from hallucinations.
-They over-rely on partial cues and generate incorrect responses.
-Recently, methods like Visual Contrastive Decoding (VCD) and Instruction Contrastive Decoding (ICD) have been proposed to mitigate hallucinations by contrasting predictions from perturbed or negatively prefixed inputs against original outputs.
-In this work, we uncover that methods like VCD and ICD fundamentally influence the model’s internal attention dynamics.
-This observation suggests that their effectiveness may not stem merely from surface-level modifications to logits but from deeper shifts in attention distribution.
-Inspired by this insight, we propose an attention-steerable contrastive decoding framework that directly intervenes in the model’s attention mechanisms to offer a more principled approach to mitigating hallucinations.
-Specifically, we introduce positive and negative steering as two complementary directions for adapting the model’s internal attention distributions.
-Rather than passively adjusting logits – as it is commonly done – our method dynamically modulates attention pathways within the contrastive decoding process.
-This enables selective enhancement or suppression of visual feature contributions in a structured manner.
-Furthermore, we propose a dynamic selection mechanism to identify text-centric heads – those that predominantly attend to text over visual features – for targeted positive steering, as well as a complementary mechanism to select the most critical visual tokens for negative steering, enabling more effective attention adjustments.
-Our experiments across multiple MLLM architectures (e.g., LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP) and diverse decoding methods (greedy search, beam search, nucleus sampling) demonstrate that our approach significantly reduces hallucinations and improves the performance on benchmarks such as POPE, CHAIR, and MMHal-Bench, while simultaneously enhancing performance on standard VQA benchmarks, including MMMU, MM-VET, ScienceQA, TextVQA, and GQA.
-ASCD: Attention-Steerable Contrastive Decoding for Reducing Hallucination in MLLM
-Yujun Wang1, Jinhe Bi2, Yunpu Ma2,3, Soeren Pirk1 1Christian-Albrechts-Universität zu Kiel (CAU) 2Ludwig Maximilian University of Munich 3Munich Center for Machine Learning
-1 Introduction
-Recent advances in large language models (LLMs) (Yang et al., 2024; Touvron et al., 2023; Abdin et al., 2024; Raffel et al., 2023; Brown et al., 2020; Devlin et al., 2019) have led to impressive results in a wide array of natural language processing tasks.
-Building on these successes, researchers have extended LLMs by visual inputs that enable multi-modal large language models (MLLMs) such as LLaVA (Liu et al., 2023b, 2024b).
-These MLLMs can handle complex tasks like image captioning Anderson et al.
-(2018), visual question answering Agrawal et al.
-(2016), and multimodal dialogue Das et al.
-(2017).
-Existing approaches (Dai et al., 2023; Liu et al., 2023b, 2024b; Zhou et al., 2024; Chen et al., 2023a; Alayrac et al., 2022; Bi et al., 2024) show remarkable potential to bridge the gap between vision and language.
-Despite these achievements, MLLMs often inherit a critical limitation from LLMs: the tendency to produce hallucinations (Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a).
+Multimodal large language models (MLLMs) frequently hallucinate by over-committing to spurious visual cues.
+Prior remedies—Visual and Instruction Contrastive Decoding (VCD, ICD)—mitigate this issue, yet the mechanism remains opaque.
+We first empirically show that their improvements systematically coincide with redistributions of cross-modal attention.
+Building on this insight, we propose Attention-Steerable Contrastive Decoding (ASCD), which directly steers the attention scores during decoding.
+ASCD combines (i) positive steering, which amplifies automatically mined text-centric heads—stable within a model and robust across domains—with (ii) negative steering, which dampens on-the-fly identified critical visual tokens.
+The method incurs negligible runtime/memory overhead and requires no additional training.
+Across five MLLM backbones and three decoding schemes, ASCD reduces hallucination on POPE, CHAIR, and MMHal-Bench by up to 38.2% while improving accuracy on standard VQA benchmarks, including MMMU, MM-VET, ScienceQA, TextVQA, and GQA.
+These results position attention steering as a simple, model-agnostic, and principled route to safer, more faithful multimodal generation.
+Introduction
+Recent advances in large language models (LLMs) [43, 40, 1, 34, 10, 18] have led to impressive results in a wide array of natural language processing tasks.
+Building on these successes, researchers have extended LLMs with visual inputs that enable multimodal large language models (MLLMs) such as LLaVA [29, 28].
+These MLLMs can handle complex tasks like image captioning [4], visual question answering [2], and multimodal dialogue [16].
+Existing approaches [15, 29, 28, 48, 12, 3, 8] show remarkable potential to bridge the gap between vision and language.
+Despite these achievements, MLLMs often inherit a critical limitation from LLMs: the tendency to produce hallucinations [20, 6, 30].
 These hallucinations arise when a model over-relies on partial or misleading cues, generating responses that are incorrect or do not correspond to the provided input.
 Figure 1: Impact of VCD and ICD on attention distribution.
-We conduct an image description task using LLaVA-1.5 on 500 randomly sampled images from the COCO dataset while monitoring the internal attention distribution within the LLM component.
-We compare the changes in attention under different settings of Visual Contrastive Decoding (VCD), Instruction Contrastive Decoding (ICD), and their combination, relative to the original LLaVA model.
-The x-axis represents different attention categories: system tokens (sys), visual tokens (vis), textual tokens (text), and output tokens (output).
-The y-axis indicates the attention difference relative to the original model.
-VCD (solid blue bars) reduces attention to visual tokens while slightly increasing attention to textual tokens, with a stronger effect as the number of noising steps increases.
-ICD (hatched bars) exhibits a similar trend, further decreasing visual attention and increasing text attention, where stronger negative prefixes (see text in the legend) result in a more pronounced shift.
-When combining VCD and ICD (dotted bars), the reduction in visual attention is further amplified, while the focus on textual tokens increases.
-These findings indicate that the effectiveness of VCD and ICD originates from underlying shifts in the model’s attention distribution rather than solely from the contrastive decoding process.
+On 500 COCO images, we measure how Visual (VCD) and Instruction (ICD) Contrastive Decoding redistribute attention in LLaVA-1.5.
+Both techniques—and their combination—lower attention on visual tokens (vis) while raising it on textual tokens (text), with stronger perturbations yielding larger shifts.
+This suggests that the reduction in hallucinations achieved by VCD and ICD is largely attributable to the attendant shifts in cross-modal attention, rather than to the logit-subtraction step alone.
 To mitigate hallucinations, two general strategies have emerged: training-phase interventions and inference-phase interventions.
-In the training phase, auxiliary supervision Chen et al.
-(2023b) or reinforcement learning Ben-Kish et al.
-(2024) can help align model outputs with factual or human-preferred references.
+In the training phase, auxiliary supervision [14] or reinforcement learning [7] can help align model outputs with factual or human-preferred references.
 However, these approaches require additional data or complex reward modeling, which may be costly or infeasible in certain scenarios.
-In contrast, inference-phase methods (Zhou et al., 2024; Zhao et al., 2024; Deng et al., 2024; Wang et al., 2024a; Leng et al., 2023) aim to correct or filter erroneous outputs without retraining.
+In contrast, inference-phase methods [48, 47, 17, 41, 24] aim to correct or filter erroneous outputs without retraining.
 Contrastive decoding is particularly appealing as it leverages negatively perturbed or prefixed inputs to steer the model away from hallucinations in a training-free manner.
-Two notable recent methods for contrastive decoding are Visual Contrastive Decoding (VCD) Leng et al.
-(2023) that perturbs an input image (e.g., via noising) to generate a “negative result” of logits, which is then subtracted from the original logits to suppress hallucinations, and second, Instruction Contrastive Decoding (ICD) Wang et al.
-(2024a) that prepends a negative prefix to the prompt (e.g., “You are a confused object detector”) to generate a signal that shifts the model’s predictions away from hallucinated content.
+Two notable recent methods for contrastive decoding are Visual Contrastive Decoding (VCD) [24] that perturbs an input image (e.g., via noising) to generate a “negative result” of logits, which is then subtracted from the original logits to suppress hallucinations, and Instruction Contrastive Decoding (ICD) [41] that prepends a negative prefix to the prompt (e.g., “You are a confused object detector”) to generate a signal that shifts the model’s predictions away from hallucinated content.
 Both methods offer a lightweight, yet effective approach to reducing hallucinations.
-However, upon closer examination, we find that these methods construct contrasting branches through surface-level modifications – either perturbing the image (VCD) or prefixing the prompt (ICD) – without explicitly addressing the underlying cause of hallucinations.
-Attention steering like OPERA and PAI (Liu et al., 2024c; Huang et al., 2024a) is also a common inference-phase remedy to reduce hallucination.
-However, PAI introduces the notion of “text inertia” – the tendency of an MLLM to keep generating text-driven content even when the image is removed – but does not articulate why steering the attention matrix is the necessary lever to overcome this inertia.
-In our experiments (Fig.
-1), we observe that both VCD and ICD consistently cause fundamental shifts in the internal attention distribution: they tend to reduce attention on visual tokens and amplify attention on textual tokens.
+However, upon closer examination, we find that these methods construct contrasting branches through surface-level modifications—either perturbing the image (VCD) or prefixing the prompt (ICD)—without explicitly addressing the underlying cause of hallucinations.
+Attention steering like OPERA, IBD and PAI [31, 50, 19] is also a common inference-phase remedy to reduce hallucinations.
+However, PAI introduces the notion of “text inertia”—the tendency of an MLLM to keep generating text-driven content even when the image is removed—but does not articulate why steering the attention matrix is the necessary lever to overcome this inertia.
+To motivate our approach, we first quantify how VCD and ICD reshape a model’s internal attention.
+As evidenced by Fig.
+1, both techniques produce a systematic reallocation of attention from visual tokens to textual tokens.
 This insight raises a natural question: why not directly steer the attention mechanism itself?
 To this end, we propose an Attention-Steerable Contrastive Decoding (ASCD) framework to manipulate attention.
-Specifically, the attention modification is integrated into a contrastive decoding pipeline to either enhance visual cues or to suppress negative signals.
+Specifically, the attention modification is integrated into a contrastive decoding pipeline to both enhance visual cues and suppress negative signals.
 We further develop a dynamic head-selection mechanism to identify “text-centric” heads that disproportionately focus on textual cues, enabling more targeted positive adjustments.
@@
 In summary, our contributions are as follows: (1) We analyze how recent contrastive decoding methods (VCD, ICD) create “negative samples” that fundamentally alter attention; (2) We propose an attention-steerable contrastive decoding method that explicitly modulates attention distributions to offer a more principled way to mitigate hallucinations in the inference phase; (3) We faithfully reproduce VCD and ICD to ensure fair comparison with prior work.
-Across three representative MLLM backbones (LLaVA-1.5 7B, LLaVA-NeXT 7B, and Phi2-SigLIP), three decoding schemes (greedy, nucleus, and beam search), and three hallucination-focused benchmarks (Rohrbach et al., 2019; Li et al., 2023b; Sun et al., 2023) (POPE, CHAIR, MMHal-Bench), our approach consistently reduces hallucinations and strengthens visual grounding.
-At the same time, it improves performance on standard VQA benchmarks (Yue et al., 2024; Yu et al., 2024; Lu et al., 2022; Singh et al., 2019; Hudson and Manning, 2019), including MMMU, MM-VET, ScienceQA, TextVQA, and GQA whereas other methods suffer from degraded performance on these benchmarks.
+Across five representative MLLM backbones (LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, LLaVA-1.5 13B and Qwen2.5-VL-Instruct), three decoding schemes (greedy, nucleus, and beam search), and three hallucination-focused benchmarks [35, 26, 39] (POPE, CHAIR, MMHal-Bench), our approach consistently reduces hallucinations and strengthens visual grounding.
+At the same time, it improves performance on standard VQA benchmarks [45, 44, 33, 37, 21], including MMMU, MM-VET, ScienceQA, TextVQA, and GQA whereas other methods suffer from degraded performance on these benchmarks.
 Figure 2: A motivating example of proactive attention steering in a visually ambiguous scenario.
-(a) shows the conversation context where the “orange” is actually tinted blue.
-(b) shows how the logits vary based on negative-steering.
-(c) shows how the logits vary based on positive-steering.
-(d) illustrates how attention-steerable contrastive decoding, which combines both negative and positive steering in a unified framework, reduce hallucinations and produce perception-driven answers.
-2 Related Work
+Top: Conversation context in which the “orange” appears blue-tinted.
+Middle: Effects of negative steering (decrease vision attention / increase text attention) and positive steering (increase vision attention / decrease text attention); ASCD contrasts the two steered logits to suppress hallucination and produce the perception-consistent answer.
+Bottom: Color-token logits change with the steering strength for visual and textual attention, corresponding to the steering above.
+Related Work
 Multimodal Large Language Models.
-Multimodal Large Language Models (MLLMs) have significantly advanced the field of artificial intelligence by integrating vision and language understanding, enabling a wide range of vision-language tasks (Dai et al., 2023; Zhu et al., 2023; Liu et al., 2024b, 2023b; Alayrac et al., 2022; Chen et al., 2023a; Zhou et al., 2024; Zhang et al., 2023a; Rong et al., 2025; Chen et al., 2025a; Yu et al., 2025; Guan et al., 2025; Huang et al., 2024c, 2025; Liu et al., 2025; Zhao et al., 2025b).
-These models typically follow a two-stage training paradigm: (1) large-scale pretraining on web-scale image-text pairs (Liu et al., 2023b; Li et al., 2023a) to learn cross-modal representations, and (2) visual instruction tuning (Liu et al., 2023a; Bi et al., 2025) on task-specific datasets to enhance multimodal instruction-following capabilities.
+Multimodal Large Language Models (MLLMs) have significantly advanced the field of artificial intelligence by integrating vision and language understanding, enabling a wide range of vision-language tasks [15, 49, 28, 29, 5, 48, 36, 11].
+These models typically follow a two-stage training paradigm: (1) large-scale pretraining on web-scale image-text pairs [29, 25] to learn cross-modal representations, and (2) visual instruction tuning [27, 9] on task-specific datasets to enhance multimodal instruction-following capabilities.
 While this paradigm has led to substantial improvements in vision-language reasoning, MLLMs still face key challenges, such as hallucination – where the model generates content that is inconsistent with the given visual input.
-(Huang et al., 2024b; Bai et al., 2024; Liu et al., 2024a).
+[20, 6, 30].
 Mitigating Hallucinations in MLLMs.
-Hallucination in MLLMs is particularly pronounced in open-ended generation tasks, where models may produce content that is not aligned with the provided visual input (Huang et al., 2024a; Jing et al., 2024; Zhang et al., 2023b).
+Hallucinations in MLLMs are particularly pronounced in open-ended generation tasks, where models may produce content that is not aligned with the provided visual input [19, 23, 46].
 Some approaches focus on the mitigation of data bias, scaling-up of vision resolution, and alignment optimization.
 Lovenia et al.
-(2024) introduce a technique that mines 95,000 negative samples by replacing original categories, attributes, or quantity information with similar but incorrect alternatives.
+[32] introduce a technique that mines 95,000 negative samples by replacing original categories, attributes, or quantity information with similar but incorrect alternatives.
 This fine-grained approach effectively enriches the contrastive signal during training, thereby enhancing the model’s robustness.
 Chen et al.
-(2024) propose InternVL, which scales the vision encoder up to 6 billion parameters and processes images with widths ranging from 1,664 to 6,144 pixels.
+[13] propose InternVL, which scales the vision encoder up to 6 billion parameters and processes images with widths ranging from 1,664 to 6,144 pixels.
 While this method improves visual detail and alignment, it requires significant computational resources for pretraining with large-scale data.
 Sun et al.
-(2023) employ Reinforcement Learning from Human Feedback (RLHF) Stiennon et al.
-(2022) to align different modalities during training.
+[39] employ Reinforcement Learning from Human Feedback (RLHF) [38] to align different modalities during training.
 This optimization strategy leads to a reduction in hallucinations by better integrating visual and textual cues.
 Bi et al.
-(2024) propose a representation steering method that effectively mitigates hallucination in multimodal models.
+[8] propose a representation steering method that effectively mitigates hallucination in multimodal models.
 Contrastive Decoding Approaches.
-Recent work has explored contrastive decoding as an effective, training-free means to mitigate hallucinations Xiao et al.
-(2025).
+Recent work has explored contrastive decoding as an effective, training-free means to mitigate hallucinations [42].
 For instance, Leng et al.
-(2023) introduced Visual Contrastive Decoding (VCD), which perturbs the input image to generate a negative logit branch that is subtracted from the original predictions, while Wang et al.
-(2024a) employs a negative prompt to steer outputs away from hallucinated content.
+[24] introduced Visual Contrastive Decoding (VCD), which perturbs the input image to generate a negative logit branch that is subtracted from the original predictions, while Wang et al.
+[41] employs a negative prompt to steer outputs away from hallucinated content.
 Huo et al.
-(2024) leverages a Context and Text-aware Token Selection (CT2S) strategy to selectively retain the most informative vision tokens in early decoder layers, thereby amplifying beneficial multimodal context and suppressing spurious hallucinations.
-3 Preliminaries
-Modern MLLMs integrate text and visual inputs based on powerful encoders that enable merging the modalities into a unified representation that is processed by a multi-layer Transformer.
-While these models enable producing coherent responses, they heavily rely on internal attention mechanisms that dictate how visual and textual cues are combined.
-As discussed in Section 3.2, subtle variations in these attention distributions can significantly impact the generated output.
+[22] leverage a Context and Text-aware Token Selection (CT2S) strategy to selectively retain the most informative vision tokens in early decoder layers, thereby amplifying beneficial multimodal context and suppressing spurious hallucinations.
+Preliminaries
+Modern MLLMs integrate text and visual inputs based on powerful encoders that enable the merging of the modalities into a unified representation that is processed by a multi-layer Transformer.
+While these models enable the production of coherent responses, they heavily rely on internal attention mechanisms that dictate how visual and textual cues are combined.
+As discussed in previous section, subtle variations in these attention distributions can significantly impact the generated output.
 This observation motivates our approach: by explicitly modulating attention, we aim to enhance visual grounding and mitigate hallucinations.
-3.1 MLLM Formulation
+MLLM Formulation
 We consider a multimodal large language model (MLLM) that processes an image $\mathbf{I}$ and a text prompt $\mathbf{x}=\{x_{1},\dots,x_{N}\}$ to generate an output sequence $\mathbf{y}=\{y_{1},\dots,y_{M}\}$ in an autoregressive manner.
@@
 where $y_{<t}$ denotes all previously generated tokens.
-Embeddings. A unified input is obtained from encoded image and embedded text:
-$\mathbf{Z}=[f_{v}(\mathbf{I});f_{t}(\mathbf{x})].$ (2)
-Transformer Architecture. The MLLM processes $\mathbf{Z}$ through $L$ Transformer layers Vaswani et al. (2023):
-$\mathbf{H}^{(l)}=\mathrm{TransformerLayer}^{(l)}\!\bigl{(}\mathbf{H}^{(l-1)}\bigr{)},\quad\mathbf{H}^{(0)}=\mathbf{Z}.$ (3)
-Output Prediction. The final hidden state $\mathbf{h}_{t}^{(L)}$ is mapped to a probability distribution over the vocabulary:
-$p_{\theta}(y_{t}\mid\mathbf{I},\mathbf{x},y_{<t})=\mathrm{Softmax}\!\Bigl{(}\mathbf{h}_{t}^{(L)}W^{P}\Bigr{)},$ (4)
-where $W^{P}$ is the output projection matrix.
-3.2 Proactive Steering of Attention
+Transformer Backbone.
+The input is processed by $L$ Transformer blocks, and each block contains $H$ attention heads. We denote the unnormalized attention score matrix of the head $h$ in layer $l$ by $\mathbf{A}^{(l)}_{h}$.
+Proactive Steering of Attention
 In Figure 1, we show how visual contrastive decoding (VCD) and instruction contrastive decoding (ICD) indirectly alter attention distributions.
 Building on this insight, we now ask: what if we explicitly steer the model’s attention?
-Figure 2 provides a motivating example, illustrating how actively modulating attention can influence the final logits distribution.
-Consider a simple query: “What is the color of the orange here?” The conversation context (Figure 2a) is based on LLaVA-1.5 7B, with a provided image in which the “orange” fruit appears to be tinted blue.
-We experiment with two distinct attention-steering scenarios: negative-steered logits (Figure 2b) and positive-steered logits (Figure 2c).
+Figure 2 provides a motivating example, illustrating how actively modulating attention can influence the final logit distribution.
+Consider a simple query: “What is the color of the orange here?” The conversation context (Figure 2) is based on LLaVA-1.5 7B, with a provided image in which the “orange” fruit appears to be tinted blue.
+We experiment with two distinct attention-steering scenarios: negative-steered logits and positive-steered logits.
 In each case, we proportionally adjust the visual or textual attention before finalizing the output distribution.
@@
 As shown in the histogram of logits, the model reduces its reliance on the visual input, causing it to fall back more heavily on the LLM’s inherent priors.
-As a result, it is more likely to generate answers that align with typical linguistic associations rather than the actual content of the image – insisting the color is “orange”.
+As a result, it is more likely to generate answers that align with typical linguistic associations rather than the actual content of the image—insisting that the color is “orange”.
 Conversely, the positive-steered branch increases attention to visual tokens or downgrades textual tokens, making the model more sensitive to the actual (albeit unexpected) color in the image.
@@
 In addition to these unidirectional adjustments, we further integrate attention steering into the contrastive decoding framework.
-Instead of using the original logits directly (as in VCD or ICD), we inject the attention-modulated logits.
+Instead of using the original logits for the positive branch directly (as in VCD or ICD), we inject the attention-modulated logits.
 Mathematically, we redefine the contrastive decoding formulation by replacing the original logits adjustment with a positively steered version:
@@
 where $p_{\theta}^{\text{pos-steered}}$ and $p_{\theta}^{\text{neg-steered}}$ represent the output logits modified by positively or negatively steered attention.
-By integrating contrastive decoding with explicit attention manipulation, our attention-steerable contrastive decoding framework (Figure 2d), sharpens the output distribution which enhances the likelihood of the correct response, while reducing the impact of competing distractors.
-Figure 3: Distribution of text-centric heads across different models and experiment settings.
+By integrating contrastive decoding with explicit attention manipulation, our attention-steerable contrastive decoding framework (Figure 2 right) sharpens the output distribution, enhances the likelihood of the correct response, while reducing the impact of competing distractors.
+Methodology
+Figure 3: The stability of text-centric head distribution.
 Each heatmap visualizes how frequently a given head occurs among the most text-focused heads.
-The panel in the center (a) show the result of LLaVA-1.5 with a generation length of 64 tokens; (b) and (c) show results of the same model with longer generation (512 tokens) and a different image set.
-Despite these changes, LLaVA-1.5 exhibits minimal JS divergence, which indicates consistent text-centric heads.
-In contrast, Phi2-SigLIP (d) and LLaVA-NeXT (e) deviate substantially from LLaVA-1.5, revealing model-specific attention biases and higher JS divergence.
-3.3 Text-centric Heads
-Previously, we have highlighted the impact of adjusting attention.
-In this section, we discuss which heads in the model are most prone to over-reliance on textual cues.
-To this end, we conduct an experiment to identify "text-centric" heads, i.e., those with disproportionately high text-to-visual attention ratios, and examine their consistency under different generation conditions and image sets.
-The experiment setup is detailed in Appendix A.
+LLaVA-1.5(a) remains stable across generation length(b) and image set(c), whereas Phi2-SigLIP(d) and LLaVA-NeXT(e) shift markedly.
+In this section, we present our attention-steerable contrastive decoding framework, which explicitly modulates the model’s attention to mitigate hallucinations.
+Our approach has two stages: (1) Text-centric Head Selection, which identifies the heads most prone to text-centric bias, and (2) Attention Steering, where we apply positive steering to text-centric heads and negative steering to a small subset of visually critical tokens.
+We then integrate these adjusted logits for generation into a contrastive decoding pipeline.
+Text-centric Heads are Model-specific
+Having established the impact of attention adjustments, we now discuss which heads in the model are most prone to over-reliance on textual cues.
+Here, we conduct an experiment to identify "text-centric" heads, i.e., those with disproportionately high text-to-visual attention ratios, and examine their consistency under different generation conditions and image sets.
+The experimental setup is detailed in the supplementary material (Appendix, Text-Centric Heads Experiment Settings).
 Results and Observations.
@@
 The panel in the center (a) corresponds to LLaVA-1.5 on $N=500$ images with a generation length of 64 tokens.
-The two heatmaps at the bottom show results of the same model but with either an increased generation length to 512 tokens (b, bottom left), or using a different set of 500 images (c, bottom right).
-Despite these changes, the distribution of top text-focus heads remains visually similar, and the small Jensen–Shannon (JS) divergences confirm that these text-centric heads are largely invariant under different sampling conditions for the same model.
-In contrast, the Phi2-SigLIP (d, top-left) and LLaVA-NeXT (e, top-right) panels deviate significantly from LLaVA-1.5 even under the same experiment settings, with higher JS divergence.
+The two heatmaps at the bottom show results of the same model but with either an increased generation length to 512 tokens (b, bottom left) or using a different set of 500 images (c, bottom right).
+Despite these changes, the distribution of top text-focused heads remains visually similar, and the small Jensen–Shannon (JS) divergences confirm that these text-centric heads are largely invariant under different sampling conditions for the same model.
+In contrast, the Phi2-SigLIP (d, top-left) and LLaVA-NeXT (e, top-right) panels deviate significantly from LLaVA-1.5 even under the same experimental settings, with higher JS divergence.
 This suggests that each model has its own unique set of heads that consistently favor textual attention over visual cues.
@@
 The consistent presence of the text-centric heads within the same model indicates that certain heads are inherently prone to focusing on textual signals rather than visual content.
-In Section 4.2 we describe how this insight can be leveraged to selectively target the problematic heads when applying our positive steering strategy.
-Rather than uniformly amplifying attention across all heads, we concentrate on those that are most responsible for text-dominant attention, thereby avoiding unnecessary modifications to heads that are well-balanced in their visual-textual focus.
-4 Methodology
+In the next subsection, we describe how this insight can be leveraged to selectively target the problematic heads when applying our positive steering strategy.
+Robustness across Data Domains.
+To further assess the robustness of model-specific text-centric heads, we repeat the profiling on extremely out-of-domain X-ray data.
+Details are provided in the supplementary material (“Extended Analysis of Text-centric Heads on Medical Data Domain”).
+The resulting heatmaps and Jensen–Shannon divergences show that, within each model, the same subset of heads remains text-centric despite the shift from COCO photographs to chest X-rays, indicating strong domain robustness.
 Figure 4: Illustration of positive and negative steering.
@@
 These selective adjustments work in tandem to reduce hallucinations and improve grounding.
-In this section, we present our attention-steerable contrastive decoding framework, which explicitly modulates the model’s attention to mitigate hallucinations.
-Our approach has two stages: (1) Text-centric Head Selection, which identifies the heads most prone to text-centric bias, and (2) Attention Steering, where we apply positive steering to text-centric heads and negative steering to a small subset of visually critical tokens.
-We then integrate these adjusted logits for generation into a contrastive decoding pipeline.
-4.1 Text-centric Head Selection
+Text-centric Head Selection
+Input: Reference image set $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$; MLLM with $L$ layers and $H$ heads per layer; desired text-centric head count $\kappa_{\textsc{tch}}$
+Output: $\mathcal{H}_{\textsc{pos}}$ // Selected text-centric heads
+Initialize Global Statistics:
+Initialize counter tensor $F\leftarrow\mathbf{0}^{L\times H}$ // Vote accumulator
+[2pt]
+Vote Accumulation Over Reference Set:
+foreach $\mathbf{I}_{i}\in\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$ do // Iterate over images
+   Run MLLM on $\mathbf{I}_{i}$ to obtain cross-modal attentions
+ foreach $(r,c)\in\{1,\dots,L\}\times\{1,\dots,H\}$ do // Loop over all heads
+    $Q_{i}(r,c)\leftarrow\dfrac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)}$
+ $\mathcal{I}_{i}\leftarrow$ indices of the top-32 $Q_{i}$ values
+ foreach $(r,c)\in\mathcal{I}_{i}$ do
+    $F(r,c)\leftarrow F(r,c)+1$
+Head Selection:
+Sort heads $(r,c)$ by $F(r,c)$ in descending order
+$\mathcal{H}_{\textsc{pos}}\leftarrow$ top $\kappa_{\textsc{tch}}$ heads
+return $\mathcal{H}_{\textsc{pos}}$
 Algorithm 1 Text-Centric Head Selection (Offline)
-1:Reference dataset $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$, MLLM with $L$ layers and $H$ heads per layer, Final text-centric head count $\kappa_{\textsc{tch}}$
-2:$\mathcal{H}_{\textsc{pos}}$ (set of selected text-centric heads)
-3:Initialize a global counter $F\in\mathbb{R}^{L\times H}$ to zeros
-4:for $i\leftarrow 1$ to $N$ do
-5:     Run MLLM on image $\mathbf{I}_{i}$ (e.g., image description)
-6:     for all head $(r,c)$ in layer-head grid do
-7:         Compute attention ratio:
-$Q_{i}(r,c)=\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)}$
-8:     end for
-9:     Identify top-32 indices of $Q_{i}$ (largest ratios) and store in $\mathcal{I}_{i}$
-10:     for all $(r,c)\in\mathcal{I}_{i}$ do
-11:         $F(r,c)\leftarrow F(r,c)+1$
-12:     end for
-13:end for
-14:Sort all heads $(r,c)$ in descending order by $F(r,c)$
-15:Select top $\kappa_{\textsc{tch}}$ heads:
-$\mathcal{H}_{\textsc{pos}}\leftarrow\text{first }\kappa_{\textsc{tch}}\text{ heads in sorted list}$
-16:return $\mathcal{H}_{\textsc{pos}}$
 As detailed in Algorithm 1, we start by identifying the most text-centric heads using a small reference dataset (e.g., 500 images) for a task (e.g., image description).
-For each sample, we compute the ratio of textual attention to visual attention (Eq.
-6 in Appendix A) and take the top 32 heads with the highest ratio.
+For each sample, we compute the ratio of textual attention to visual attention and take the top 32 heads with the highest ratio.
 We accumulate these counts over all samples, then choose the top $\kappa_{\textsc{tch}}$ heads as “text-centric”.
-This step is motivated by our finding (Section 3.3) that certain heads consistently favor textual content over visual cues.
-Algorithm 2 Attention-Steerable Contrastive Decoding
-1:Image $\mathbf{I}$, Text-centric heads $\mathcal{H}_{\textsc{pos}}$ (from Algorithm 1), Critical vis-token count $\kappa_{\textsc{vis}}$, Steer strengths $\alpha_{\textsc{pos}},\alpha_{\textsc{neg}}$, Contrastive weight $\alpha$, Truncation threshold $\beta$, MLLM with $L$ layers and $H$ heads per layer
-2:$p_{\theta}^{\text{final}}$ (final logits from ASCD)
-3:Step 1: Forward Pass with Positive Steering
-4:for $l\leftarrow 1$ to $L$ do
-5:     for $h\leftarrow 1$ to $H$ do
-6:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
-7:         if $(l,h)\in\mathcal{H}_{\textsc{pos}}$ then
-8:              $\mathbf{A}^{(l)}_{h}\;\leftarrow\;\mathbf{A}^{(l)}_{h}\;+\;\alpha_{\textsc{pos}}\,|\mathbf{A}^{(l)}_{h}|$
-9:         end if
-10:     end for
-11:     Normalize $\mathbf{A}^{(l)}$ and continue
-12:end for
-13:Obtain logits $p_{\theta}^{\text{pos-steered}}$
-14:Step 2: Forward Pass with Negative Steering
-15:for $l\leftarrow 1$ to $L$ do
-16:     for $h\leftarrow 1$ to $H$ do
-17:         Compute attention matrix $\mathbf{A}^{(l)}_{h}$
-18:         Identify top-$\kappa_{\textsc{vis}}$ critical visual tokens $\mathcal{V}$
-19:         for all $v\in\mathcal{V}$ do
-20:              $\mathbf{A}^{(l)}_{h}(v)\;\leftarrow\;\mathbf{A}^{(l)}_{h}(v)\;-\;\alpha_{\textsc{neg}}\,|\mathbf{A}^{(l)}_{h}(v)|$
-21:         end for
-22:     end for
-23:     Normalize $\mathbf{A}^{(l)}$ and continue
-24:end for
-25:Obtain logits $p_{\theta}^{\text{neg-steered}}$
-26:Step 3: Contrastive Decoding with Truncation
-27:$p_{\theta}^{\text{raw}}\;\leftarrow\;(1+\alpha)\,p_{\theta}^{\text{pos-steered}}\;-\;\alpha\,p_{\theta}^{\text{neg-steered}}$
-[…]
-29:$p_{\theta}^{\text{final}}\;\leftarrow\;p_{\theta}^{\text{raw}}.\text{masked\_fill}(p_{\theta}^{\text{pos-steered}}<\text{cutoff},-\infty)$
-30:return $p_{\theta}^{\text{final}}$
-4.2 Attention Steering
+This step is motivated by our previous finding that certain heads consistently favor textual content over visual cues.
+Input: Image $\mathbf{I}$; text-centric heads $\mathcal{H}_{\textsc{pos}}$; critical visual-token count $\kappa_{\textsc{vis}}$; steer strengths $\alpha_{\textsc{pos}},\alpha_{\textsc{neg}}$; contrastive weight $\alpha$; truncation threshold $\beta$; MLLM with $L$ layers and $H$ heads
+Output: $p_{\theta}^{\text{final}}$
+1. Positive Steering Pass:
+for $l\leftarrow 1$ to $L$ do
+ for $h\leftarrow 1$ to $H$ do
+      Compute attention matrix $\mathbf{A}^{(l)}_{h}$
+    if $(l,h)\in\mathcal{H}_{\textsc{pos}}$ then
+       $\mathbf{A}^{(l)}_{h}\leftarrow\mathbf{A}^{(l)}_{h}+\alpha_{\textsc{pos}}\lvert\mathbf{A}^{(l)}_{h}\rvert$
+  Normalize $\mathbf{A}^{(l)}$ and continue
+Obtain logits $p_{\theta}^{\text{pos}}$
+[4pt]
+2. Negative Steering Pass:
+for $l\leftarrow 1$ to $L$ do
+ for $h\leftarrow 1$ to $H$ do
+      Compute attention matrix $\mathbf{A}^{(l)}_{h}$
+      Identify top-$\kappa_{\textsc{vis}}$ critical visual tokens $\mathcal{V_{\text{crit}}}$
+    foreach $v\in\mathcal{V_{\text{crit}}}$ do
+       $\mathbf{A}^{(l)}_{h}(v)\leftarrow\mathbf{A}^{(l)}_{h}(v)-\alpha_{\textsc{neg}}\lvert\mathbf{A}^{(l)}_{h}(v)\rvert$
+  Normalize $\mathbf{A}^{(l)}$ and continue
+Obtain logits $p_{\theta}^{\text{neg}}$
+[4pt]
+3. Contrastive Decoding and Truncation:
+$p_{\theta}^{\text{raw}}\leftarrow(1+\alpha)\,p_{\theta}^{\text{pos}}-\alpha\,p_{\theta}^{\text{neg}}$
+$\text{cutoff}\leftarrow\log(\beta)+\max(p_{\theta}^{\text{raw}})$
+$p_{\theta}^{\text{final}}\leftarrow p_{\theta}^{\text{raw}}.\text{masked\_fill}(p_{\theta}^{\text{pos}}<\text{cutoff},-\infty)$
+return $p_{\theta}^{\text{final}}$
+Algorithm 2 Attention-Steerable Contrastive Decoding (ASCD)
+Attention Steering
 Text-centric Head Awareness and Critical Visual Token Selection.
@@
 Figure 5a highlights how targeted steering in text-centric heads improves the positive steering effectiveness.
-Simultaneously, we apply negative steering to the top $\kappa_{\textsc{vis}}$ most critical visual tokens – those receiving the highest aggregate attention across heads – reducing their attention scores by $\alpha_{\textsc{neg}}$ across all heads.
+Simultaneously, we perform negative steering on the $\kappa_{\textsc{vis}}$ visual tokens that draw the most attention—those with the highest head-averaged score.
+For the current query position, we define
+$s(v)=\frac{1}{H}\sum_{h=1}^{H}A^{(L)}_{h}(v),\qquad v\in\{1,\dots,V\}.$ (3)
+The $\kappa_{\textsc{vis}}$ tokens with the largest $s(v)$ values form the critical set $\mathcal{V}_{\text{crit}}$; every attention entry to any $v\in\mathcal{V}_{\text{crit}}$ is then down-scaled in all heads by $\alpha_{\textsc{neg}}\lvert A^{(L)}_{h}(v)\rvert$.
 Through this strategy, we deliberately obscure only the most pivotal cues – this targeted suppression is sufficient to induce a strong hallucination effect in the negative branch, leading to improved contrastive decoding compared to a blanket suppression of all visual tokens.
@@
 Figure 5: Comparative effectiveness of selective attention steering.
-(a): Positive steering applied only to text-centric heads (32 heads with the highest text-to-visual ratio) outperforms random or blanket head selection across various decoding strategies (Greedy, Nucleus, Beam), leading to higher POPE Accuracy and F1.
-(b): Negative steering focused on a small subset of critical visual tokens, integrated with contrastive decoding, significantly reduces CHAIR metrics (less hallucination) and boosts POPE metrics compared to randomly suppressing visual tokens of the same number.
-These results validate that targeted attention modulation on text-centric heads (for positive steering) and critical visual tokens (for negative steering) yields stronger hallucination mitigation and more grounded responses.
+(a): Positive steering applied only to text-centric heads outperforms random or blanket head selection across various decoding strategies.
+(b): Negative steering focused on a small subset of critical visual tokens, integrated with contrastive decoding, significantly reduces CHAIR metrics (less hallucination) and boosts POPE scores compared to randomly suppressing visual tokens of the same number.
 Integration with Contrastive Decoding with Truncation.
-Building on the attention-steering process, we first obtain two output distributions: $p_{\theta}^{\text{pos-steered}}$ from the positively steered branch and $p_{\theta}^{\text{neg-steered}}$ from the negatively steered branch.
-We then combine these into contrastive decoding with a truncation mechanism, as detailed in the Step 3 of Algorithm 2.
-This process not only reinforces visually grounded predictions but also effectively mitigates the influence of spurious textual biases.
-5 Experiments
+We first compute two output distributions: $p_{\theta}^{\text{pos}}$ from the positively steered branch and $p_{\theta}^{\text{neg}}$ from the negatively steered branch.
+Step 3 of Algorithm 2 then fuses them through a contrastive-decoding rule with truncation, producing the final logits.
+This fusion amplifies visually grounded evidence while suppressing spurious text-only cues, thereby reducing hallucinations.
+Experiments
+Model Decoding Method CHAIRs (↓) CHAIRi (↓) POPE-Acc (↑) POPE-F1 (↑)
+LLaVA-
+1.5 7B
+greedy Orig 53.2 13.5 85.37 84.06
+VCD 56.8 15.2 84.27 83.35
+ICD 52.8 13.2 83.07 80.64
+PAI - - 85.82 85.79
+ASCD 35.6 (33.1%) 8.6 (36.3%) 86.53 86.25
+nucleus Orig 59.0 17.4 83.03 81.57
+VCD 59.8 16.6 83.31 82.30
+ICD 57.4 15.6 82.13 79.62
+PAI - - 81.72 82.87
+ASCD 43.6 (26.1%) 11.3 (35.1%) 85.75 85.07
+beam Orig 54.8 15.3 85.40 84.10
+VCD 58.8 16.4 84.27 83.30
+ICD 52.6 13.9 83.04 80.59
+PAI - - 86.33 85.89
+ASCD 40.8 (25.5%) 10.1 (34.0%) 86.52 86.24
+LLaVA-
+NeXT 7B
+greedy Orig 31.6 7.5 83.93 81.89
+VCD 37.2 9.7 84.86 83.28
+ICD 32.8 8.4 84.44 82.70
+ASCD 21.8 (31.0%) 7.0 (6.7%) 84.85 83.40
+nucleus Orig 30.4 8.0 81.74 79.61
+VCD 40.4 10.4 83.55 81.95
+ICD 39.4 9.9 83.67 81.98
+ASCD 21.2 (30.3%) 6.7 (16.3%) 84.69 83.09
+beam Orig 34.0 8.5 84.11 82.14
+VCD 36.6 9.1 84.66 83.03
+ICD 31.8 7.6 84.48 82.75
+ASCD 21.0 (38.2%) 6.5 (23.5%) 84.91 83.48
+Phi2-
+SigLIP
+greedy Orig 29.0 6.9 87.19 86.16
+VCD 39.4 9.6 86.22 85.53
+ICD 33.4 7.7 85.83 84.58
+ASCD 21.8 (24.8%) 5.4 (21.7%) 87.81 86.90
+nucleus Orig 36.0 9.8 85.51 84.44
+VCD 36.0 8.1 85.60 84.86
+ICD 37.0 9.4 84.63 83.35
+ASCD 26.0 (27.8%) 8.0 (18.4%) 87.45 86.46
+beam Orig 30.4 6.9 87.19 86.16
+VCD 36.0 8.4 86.30 85.64
+ICD 31.0 7.0 85.83 84.58
+ASCD 24.6 (19.1%) 5.7 (17.4%) 87.81 86.90
+LLaVA-
+1.5 13B
+greedy Orig 51.2 12.6 85.52 84.12
+ASCD 33.0(35.5%) 8.0(36.5%) 87.78 87.40
+nucleus Orig 51.4 14.5 83.95 82.54
+ASCD 35.7(30.5%) 9.5(34.5%) 87.19 86.46
+Qwen-
+2.5-VL-
+Instruct
+greedy Orig 31.2 7.5 87.72 86.64
+ASCD 24.4(21.8%) 6.4(14.7%) 88.91 88.27
+nucleus Orig 34.8 8.6 87.34 86.14
+ASCD 26.9(22.7%) 6.3(26.7%) 88.45 87.68
+Table 1: CHAIR and POPE Evaluation Results.
+Lower CHAIRs and CHAIRi values indicate better performance in reducing hallucinations.
+POPE performance is reported as the mean accuracy and F1 score.
+The best values for each metric within a model-decoding combination are highlighted in bold.
+If ASCD ranks second, the best is bold while the ASCD score is underlined.
 To evaluate the effectiveness of our attention-steerable contrastive decoding framework in mitigating hallucinations in MLLMs, we conduct a range of experiments.
-This includes three diverse benchmarks – CHAIR, POPE, and MMHal-Bench – each designed to assess different aspects of object hallucinations.
-To ensure the broad applicability and robustness of our approach, we also test it on three representative models – LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, and employ three different decoding strategies: greedy search, nucleus sampling, and beam search.
-Details of the experimental settings are provided in Appendix B.
-Furthermore, we evaluate performance on standard VQA benchmarks including MMMU, MM-VET, ScienceQA, TextVQA, and GQA to verify that the proposed method preserves – rather than diminishes – the model’s original visual understanding.
+This includes three diverse benchmarks—CHAIR, POPE, and MMHal-Bench—each designed to assess different aspects of object hallucinations.
+To ensure the broad applicability and robustness of our approach, we test it on three representative models—LLaVA-1.5 7B, LLaVA-NeXT 7B, Phi2-SigLIP, and employ three different decoding strategies: greedy search, nucleus sampling, and beam search.
+Details of the experimental settings are provided in the supplementary material (Appendix, Evaluation Settings).
+Furthermore, we evaluate performance on standard VQA benchmarks including MMMU, MM-VET, ScienceQA, TextVQA, and GQA to verify that the proposed method preserves—rather than diminishes—the model’s original visual understanding.
+In addition, an extended evaluation on the larger LLaVA-1.5 13B and the modern Qwen2.5-VL-Instruct is conducted to verify scalability and architectural generality on a subset of benchmarks.
 Figure 6: Radar charts of MMHal-Bench results. Each axis represents a different evaluation dimension in MMHal-Bench, and a larger enclosed area indicates better overall performance.
-Model Decoding Method CHAIRs (↓) CHAIRi (↓)
-LLaVA-1.5 7B greedy Orig 53.2 13.5
-VCD 56.8 15.2
-ICD 52.8 13.2
-ASCD 35.6 8.6
-nucleus Orig 59.0 17.4
-VCD 59.8 16.6
-ICD 57.4 15.6
-ASCD 43.6 11.3
-beam Orig 54.8 15.3
-VCD 58.8 16.4
-ICD 52.6 13.9
-ASCD 40.8 10.1
-LLaVA-NeXT 7B greedy Orig 31.6 7.5
-VCD 37.2 9.7
-ICD 32.8 8.4
-ASCD 21.8 7.0
-nucleus Orig 30.4 8.0
-VCD 40.4 10.4
-ICD 39.4 9.9
-ASCD 21.2 6.7
-beam Orig 34.0 8.5
-VCD 36.6 9.1
-ICD 31.8 7.6
-ASCD 21.0 6.5
-Phi2-SigLIP greedy Orig 29.0 6.9
-VCD 39.4 9.6
-ICD 33.4 7.7
-ASCD 21.8 5.4
-nucleus Orig 36.0 9.8
-VCD 36.0 8.1
-ICD 37.0 9.4
-ASCD 26.0 8.0
-beam Orig 30.4 6.9
-VCD 36.0 8.4
-ICD 31.0 7.0
-ASCD 24.6 5.7
-Table 1: CHAIR Evaluation Results.
-Lower CHAIRs and CHAIRi values indicate better performance in reducing hallucination.
-The best values for each metric within a model-decoding combination are highlighted in bold.
 It is important to note that current benchmarks for evaluating multimodal models are highly variable.
@@
 To address these issues, we faithfully reproduced both VCD and ICD using the parameters specified in their original papers and repositories, ensuring that our evaluations are conducted under consistent conditions.
-This allows for a more reliable comparison between our method and existing approaches.
-Model Decoding Method Popular Acc Popular F1 Random Acc Random F1 Adv. Acc Adv. F1 Avg. Acc Avg. F1
+POPE and CHAIR.
+Table 1 summarizes both caption-level (CHAIR) and VQA-style (POPE) results.
+Across every backbone and decoding scheme, ASCD produces the lowest CHAIR scores and the highest POPE accuracy/F1, outperforming Orig, VCD, and ICD.
+These gains hold for all three prompt types (random, popular, adversarial; see Appendix “Detailed POPE Results”) and remain stable when scaling from 7 B to 13 B or switching to the Qwen-VL architecture, indicating that attention steering mitigates object-level hallucinations in a model- and domain-robust manner.
+MMHal-Bench.
+Figure 6 illustrates the radar charts of MMHal-Bench results for LLaVA-1.5 7B under greedy and nucleus decoding.
+Each axis represents a sub-dimension of the benchmark, and a larger area signifies better overall performance.
+ASCD exhibits the largest enclosed area, outperforming baseline, VCD, and ICD in most dimensions.
+Standard VQA Benchmarks.
+To verify that ASCD does not sacrifice a model’s general visual-question-answering ability, it’s evaluated on five widely-used VQA datasets.
+Across all three representative backbones and all decoding strategies, ASCD either matches or surpasses the original model on every dataset, while VCD and ICD consistently degrade performance as shown in Table 2.
+Summary. Our experiments confirm that ASCD effectively reduces hallucinations and improves alignment with visual content, regardless of the model or decoding strategy employed.
+Additional Experiments.
+The supplementary material presents a runtime/memory profile and a comprehensive hyper-parameter ablation.
+Together, these studies confirm that ASCD adds only marginal computational overhead and remains effective across a broad range of settings, making it practical to deploy with minimal tuning effort.
+Benchmark Orig VCD ICD ASCD
+MM-VET (↑) 31.2 30.3 33.2 33.2
+ScienceQA (↑) 67.55 67.55 67.32 69.51
+GQA (↑) 61.28 59.38 59.99 61.27
+TextVQA (↑) 57.82 55.07 57.66 57.91
+MMMU (↑) 0.342 0.333 0.360 0.348
+Table 2: Scores on five VQA-style benchmarks. The best score in each row is bold. If ASCD ranks second, the best is bold while the ASCD score is underlined.
+Conclusion
+We have shown that existing contrastive methods (e.g., VCD and ICD) inadvertently shift the internal attention distribution in MLLMs, prompting us to investigate a more direct and principled way to modulate attention.
+We propose an attention-steerable contrastive decoding framework that positively steers text-centric heads while negatively steering only the most critical visual tokens.
+Our method consistently reduces hallucinations on CHAIR, POPE, and MMHal-Bench, outperforming both baseline and previous contrastive approaches with improved and uncompromised general VQA capability.
+By targeting precisely those heads and tokens, we effectively mitigate spurious textual biases while preserving essential visual context.
+Appendix
+Appendix A Text-centric Heads Experiment Settings
+The following setup applies to "Subsection: Text-centric Heads are Model-specific".
+We select $N=500$ images $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$ (from COCO) and run an MLLM (LLaVA-1.5) in an image description task.
+During each generation, we track the ratio of textual attention to visual attention for every head:
+$Q_{i}\in\mathbb{R}^{R\times C},\quad Q_{i}(r,c)\;=\;\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)},$ (4)
+where $r$ and $c$ index each head (for instance, $R=C=32$). We then identify the top-$k$ heads with the highest ratio values and mark them in a binary mask:
+$M_{i}(r,c)\;=\;\begin{cases}1,&\text{if }(r,c)\in\text{top-$k$ indices of }Q_{i},\\ 0,&\text{otherwise.}\end{cases}$ (5)
+Finally, we aggregate these masks across all $N$ images:
+$F\;=\;\sum_{i=1}^{N}M_{i},$ (6)
+so that $F(r,c)$ records how frequently head $(r,c)$ appears among the most text-focused heads.
+Appendix B Extended Analysis of Text-centric Heads on Medical Data Domain
+Goal.
+The heads identified as text-centric were profiled on a COCO reference set containing natural photographs.
+This is already a very general and ideal setting for selecting text-centric heads.
+However, we would like to verify that this selection is not tied to the training domain but even in an extremely OOD setting, we repeat the profiling procedure on a visually dissimilar corpus of 500 NIH ChestX-ray14 images.
+Protocol.
+For every backbone we rerun the counting algorithm on the X-ray set, obtain a new head-frequency tensor $F_{\text{X-ray}}$, and compute the Jensen–Shannon divergence (JSD) between the COCO and X-ray distributions.
+For comparison we also report inter-model JSD values within each domain.
+Intra-model JS Divergence (COCO vs. X-ray)
+LLaVA-1.5 LLaVA-NeXT Phi-2-SigLIP
+JSD 0.101 0.094 0.092
+Inter-model JS Divergence on COCO
+L-1.5 vs. L-NeXT L-NeXT vs. P-S L-1.5 vs. P-S
+JSD 0.619 0.611 0.624
+Inter-model JS Divergence on X-ray
+L-1.5 vs. L-NeXT L-NeXT vs. P-S L-1.5 vs. P-S
+JSD 0.604 0.572 0.620
+Table 3: Jensen–Shannon divergence between text-centric-head distributions. A value below 0.1 is widely regarded as negligible drift. Abbreviations: L = LLaVA, P-S = Phi-2-SigLIP.
+Result.
+As shown in Table 3, all three backbones exhibit intra-model JSD $\leq 0.11$, signalling that the same set of heads remains text-centric even when the reference data shift significantly from COCO to chest X-rays.
+By contrast, swapping architectures while keeping the domain fixed yields JSD ${\sim}0.60$, confirming that the identity of text-centric heads is governed by model internals rather than by the reference dataset.
+Appendix C Extended Evaluation: Model Scaling and Modern Architecture
+Further Evaluation
+We further benchmark ASCD on (i) LLaVA-1.5-13B, to test scaling effects, and (ii) Qwen2.5-VL-Instruct, one of the latest open‐source vision–language models.
+Table 4 reports POPE accuracy/F1 under greedy and nucleus decoding.
+Across both backbones and decoding schemes, ASCD yields consistent gains, demonstrating that our attention-steering approach scales beyond 3B and 7B models and transfers across different architectures.
+Model Dec. Method Popular Random Adversarial Average
+Acc
+F1
+Acc
+F1
+Acc
+F1
+Acc
+F1
+LLaVA-
+1.5 13B
+greedy Orig 0.860 0.844 0.866 0.853 0.840 0.826 0.855 0.841
+ASCD 0.883 0.878 0.899 0.896 0.848 0.847 0.877 0.874
+nucleus Orig 0.842 0.827 0.848 0.835 0.826 0.814 0.839 0.825
+ASCD 0.876 0.867 0.887 0.880 0.851 0.845 0.871 0.864
+Qwen2.5-
+VL-Instruct
+greedy Orig 0.878 0.865 0.886 0.876 0.868 0.857 0.877 0.866
+ASCD 0.891 0.883 0.902 0.896 0.875 0.868 0.889 0.882
+nucleus Orig 0.876 0.864 0.880 0.870 0.862 0.850 0.873 0.861
+ASCD 0.889 0.881 0.895 0.888 0.869 0.860 0.884 0.876
+Table 4: POPE benchmark results on a larger (LLaVA-1.5 13B) and a novel (Qwen2.5-VL-Instruct) backbone.
+Bold numbers mark improvements of ASCD over the corresponding original model under the same decoding strategy; gray rows highlight ASCD.
+Visualization of Text-centric Heads Distribution of LLaVA-1.5 13B and Qwen2.5-VL-Instruct
+Figure 7 visualizes the head-frequency heatmap for LLaVA-1.5 13B and Qwen2.5-VL-Instruct.
+Both models reproduce the sparse, high-contrast pattern already observed for the previous backbones: a small set of heads appear in dark colors, indicating that they are marked text-centric in the vast majority of reference samples, while the remainder receive little or no votes.
+Figure 7: Text-centric-head frequency maps.
+Each pixel denotes the probability that a given layer–head is selected as text-centric when profiled on 500 COCO images.
+Darker colours highlight a handful of heads that dominate textual attention, revealing the same long-tailed, head-level anisotropy seen in previous models.
+Appendix D Detailed POPE Results
+Table 5 presents the full POPE breakdown by different type—random, popular, and adversarial—for each backbone and decoding scheme.
+We report accuracy and F1.
+Across all settings, ASCD consistently matches or surpasses the baselines, with the largest gains typically appearing on the random split.
+Model Dec. Method Popular Random Adversarial Average
+Acc
+F1
+Acc
+F1
+Acc
+F1
+Acc
+F1
 LLaVA-1.5 7B greedy Orig 85.83 84.35 86.67 85.51 83.60 82.31 85.37 84.06
@@
 If our ASCD achieves the second-best result, it is additionally marked with an underline.
-CHAIR.
-Table 1 shows the CHAIR metrics (CHAIRs and CHAIRi), which measure object hallucination in image captioning.
-Across all models and decoding strategies, ASCD consistently achieves lower CHAIR values than Orig, VCD, or ICD, which illustrates ASCD’s effectiveness at mitigating object-level hallucinations.
-POPE.
-Table 2 reports the accuracy and F1 scores under the POPE evaluation, which probes object presence with random, popular, and adversarial queries.
-Higher values indicate fewer hallucinations.
-Again, ASCD achieves the best or near-best performance in all cases.
-These gains persist across different model architectures, suggesting that attention steering is robust to model size and design variations.
-MMHal-Bench.
-Figure 7 illustrates the radar charts of MMHal-Bench results for LLaVA-1.5 7B under greedy and nucleus decoding.
-Each axis represents a sub-dimension of the benchmark, and a larger area signifies better overall performance.
-ASCD exhibits the largest enclosed area, outperforming baseline, VCD, and ICD in most dimensions.
-This improvement aligns with the CHAIR and POPE findings, underscoring the benefit of selectively steering attention to reduce hallucinations.
-Standard VQA Benchmarks.
-To verify that ASCD does not sacrifice a model’s general visual-question-answering ability, it’s evaluated on five widely-used VQA datasets.
-Across all three backbones and all decoding strategies, ASCD either matches or surpasses the original model on every dataset, while VCD and ICD consistently degrade performance as shown in the Appendix C.
-Summary. Our experiments confirm that ASCD effectively reduces hallucinations and improves alignment with visual content, regardless of the model or decoding strategy employed.
-6 Conclusion
-We have shown that existing contrastive methods (e.g.
-VCD and ICD) inadvertently shift the internal attention distribution in MLLM, prompting us to investigate a more direct and principled way to modulate attention.
-We proposed an attention-steerable contrastive decoding framework that positively steers text-centric heads while negatively steering only the most critical visual tokens.
-Our method consistently reduces hallucinations on CHAIR, POPE, and MMHal-Bench, outperforming both baseline and previous contrastive approaches with improved and uncompromised general VQA capability.
-By targeting precisely those heads and tokens, we effectively mitigate spurious textual biases while preserving essential visual context.
-Limitations
-A key limitation of ASCD is its incompatibility with FlashAttention.
-Because our method must dynamically modify the attention matrix at inference time, it cannot make use of the fused kernels, leading to higher memory consumption and slower decoding.
-A promising workaround is to distill the steering signal into the model during training: we can add an auxiliary loss – e.g., a KL-divergence term – that drives the native attention distribution to approximate the ASCD target distribution.
-If successful, the model would internalise the hallucination-mitigation behaviour, removing the need for on-the-fly edits and restoring FlashAttention speed-ups.
-We regard training-time attention regularization as a promising direction: it could distill the hallucination-robust behaviour discovered by training-free, attention-modified methods into the model itself, so that at inference the model retains this robustness while fully benefiting from FlashAttention’s speed and memory efficiency.
-Appendix A Text-centric Heads Experiment Settings
-The following setup applies to Section 3.3.
-We select $N=500$ images $\{\mathbf{I}_{1},\dots,\mathbf{I}_{N}\}$ (from COCO) and run an MLLM (LLaVA-1.5) in an image description task.
-During each generation, we track the ratio of textual attention to visual attention for every head:
-$Q_{i}\in\mathbb{R}^{R\times C},\quad Q_{i}(r,c)\;=\;\frac{\text{textAttn}(r,c)}{\text{visAttn}(r,c)},$ (6)
-where $r$ and $c$ index each head (for instance, $R=C=32$). We then identify the top-$k$ heads with the highest ratio values and mark them in a binary mask:
-$M_{i}(r,c)\;=\;\begin{cases}1,&\text{if }(r,c)\in\text{top-$k$ indices of }Q_{i},\\ 0,&\text{otherwise.}\end{cases}$ (7)
-Finally, we aggregate these masks across all $N$ images:
-$F\;=\;\sum_{i=1}^{N}M_{i},$ (8)
-so that $F(r,c)$ records how frequently head $(r,c)$ appears among the most text-focused heads.
-Appendix B Evaluation Settings
-B.1 Baseline Models and Decoding Methods
+Appendix E Evaluation Settings
+Baseline Models and Decoding Methods
 We evaluate our proposed approach using three representative models and multiple decoding strategies to demonstrate its broad applicability and robustness.
@@
 Phi2-SigLIP leverages the Phi-2 backbone and a SigLIP-based vision tower, and is trained on the ShareGPT4V dataset, offering a compact alternative with competitive capabilities.
+LLaVA-1.5 13B upsizes LLaVA-1.5 to 13B parameters—doubling language capacity while retaining the original architecture—so it serves as a clean test bed for model-scaling effects.
+Qwen2.5-VL-Instruct pairs the recent Qwen2.5-7B language model with a revamped vision encoder and multi-stage instruction tuning, representing a modern open-source architecture that achieves state-of-the-art performance out of the box.
 To assess the reliability and generalizability, we experiment with three decoding strategies: greedy search, nucleus sampling, and beam search.
-Greedy decoding yields deterministic outputs, while nucleus sampling and beam search enable for more diverse generation.
-B.2 Datasets
-With the widespread adoption of deep learning across computer vision and natural language processing Lu et al.
-(2023); Chen et al.
-(2025b, a); Cui et al.
-(2025); Yuan et al.
-(2025); Lu et al.
-(2024a, b); Zhao et al.
-(2025a, c); Xu et al.
-(2024, 2025), large language models (LLMs) have rapidly become foundational tools in AI systems Wang et al.
-(2024b); Ma et al.
-(2025); Wang et al.
-(2025b, a); Zeng et al.
-(2024); Xin et al.
-(2025); Du et al.
-(2021, 2025); Zhang et al.
-(2023c).
+Greedy decoding yields deterministic outputs, while nucleus sampling and beam search enable more diverse generation.
+Datasets
+With the widespread adoption of deep learning across computer vision and natural language processing, large language models (LLMs) have rapidly become foundational tools in AI systems.
 However, despite their remarkable capabilities, MLLMs remain susceptible to hallucination—generating descriptions inconsistent with the input image.
@@
 The Massive Multi-discipline Multimodal Understanding and Reasoning (MMMU) benchmark gauges college-level, expert-style reasoning across six broad disciplines—Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering.
-B.3 Hyperparameters
-The experiments are conducted with the following hyperparameter settings to obtain the best result:
-$\kappa_{\textsc{vis}}=0.1$, $\alpha_{\textsc{neg}}=1.0$
-LLaVA-1.5: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.6$, $\alpha=1.0$, and $\beta=0.5$.
-LLaVA-NeXT: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.7$, $\alpha=1.0$, and $\beta=0.7$.
-Phi2-SigLIP: $\kappa_{\textsc{tch}}=128$, $\alpha_{\textsc{pos}}=0.8$, $\alpha=0.5$, and $\beta=0.1$.
-Appendix C Standard VQA Capability
-Our approach substantially curbs hallucinations through explicit attention steering while simultaneously enhancing performance on standard VQA tasks, a trade-off that competing methods typically fail to avoid.
-C.1 MM-VET
-Method MM-VET Score
-Orig 31.2
-VCD 30.3
-ICD 33.2
-ASCD 33.2
-Table 3: MM-VET scores for different methods.
-C.2 ScienceQA
-Method SQA Score
-Orig 67.55
-VCD 67.55
-ICD 67.32
-ASCD 69.51
-Table 4: SQA scores for different methods.
-C.3 GQA
-Method GQA Score
-Orig 61.28
-VCD 59.38
-ICD 59.99
-ASCD 61.27
-Table 5: GQA scores for different methods.
-C.4 TextVQA
-Method TextVQA Score
-Orig 57.82
-VCD 55.07
-ICD 57.66
-ASCD 57.91
-Table 6: TextVQA scores for different methods.
-C.5 MMMU
-Method MMMU Score
-Orig 0.342
-VCD 0.333
-ICD 0.360
-ASCD 0.348
-Table 7: MMMU scores for different methods.
-Appendix D Visualization Example
-Figure 7: Oringinal LLaVA-1.5 drifts completely, describing a riverside sidewalk, parked cars, and a stroller that do not exist in the image.
+Hyperparameters
+The experiments are conducted with the following hyperparameter settings to obtain the best result on CHAIR:
+$\kappa_{\textsc{vis}}=0.1$, $\alpha_{\textsc{neg}}=1.0$, $\alpha=1.0$, and $\beta=0.1$.
+LLaVA-1.5: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.6$.
+LLaVA-NeXT: $\kappa_{\textsc{tch}}=32$, $\alpha_{\textsc{pos}}=0.7$.
+Phi2-SigLIP: $\kappa_{\textsc{tch}}=128$, $\alpha_{\textsc{pos}}=0.8$.
+Appendix F Visualization Example
+Figure 8 compares the image–caption pairs produced by the original LLaVA-1.5, its VCD and ICD variants, and our ASCD (all with greedy decoding).
+Figure 8: Original LLaVA-1.5 drifts completely, describing a riverside sidewalk, parked cars, and a stroller that do not exist in the image.
 VCD and ICD correctly identify the beach and ATV but still hallucinate extra objects such as a surfboard and a boat (highlighted in yellow).
 Our ASCD caption mentions only entities that are actually present and omits non-existent items, illustrating the improved visual grounding achieved through attention-steerable contrastive decoding.
+$\alpha_{\text{pos}}$ Acc F1
+Orig 85.4 84.1
+0.1 86.0 85.1
+0.2 85.9 85.1
+0.3 85.8 85.0
+0.4 86.0 85.3
+0.5 86.3 85.9
+0.6 86.6 86.3
+0.7 85.2 84.2
+$\alpha$ Acc F1
+Orig 85.4 84.1
+0.25 86.1 85.3
+0.50 86.4 85.9
+0.75 86.7 86.4
+1.00 86.6 86.3
+1.25 86.5 86.3
+1.50 86.6 86.4
+1.75 86.6 86.4
+2.00 86.6 86.4
+$\beta$ Acc F1
+Orig 85.4 84.1
+0.2 86.1 86.0
+0.4 86.4 86.2
+0.5 86.6 86.3
+0.6 86.6 86.2
+0.8 86.3 85.7
+1.0 85.7 84.7
+$\kappa_{\text{TCH}}$ Acc F1
+Orig 85.4 84.1
+16 86.3 85.7
+32 86.6 86.3
+64 86.6 86.3
+128 86.5 86.2
+256 86.3 85.7
+512 86.0 85.4
+Table 6: Ablation on key hyper-parameters (LLaVA-1.5 7B on POPE). Each block varies one hyper-parameter while fixing the others; bold marks the best setting per block.
+Appendix G Runtime and Memory Overhead
+Motivation.
+Attention steering operates on internal matrices at generation time and could in principle introduce non-trivial latency or memory footprints.
+To verify that this is not a practical bottleneck, we benchmark ASCD against the two mainstream contrastive-decoding baselines—VCD and ICD—using the original LLaVA-1.5-7B model.
+Setup.
+For a single A100-80G we measure (i) wall-clock latency to produce an answer for a $224\times 224$ image–question pair and (ii) peak GPU memory, under three attention back-ends: the default PyTorch eager kernel, fused SDPA, and FlashAttention 2.
+Metric Orig VCD ICD ASCD
+Latency — eager (s) 0.09 0.17 0.17 0.16
+Peak VRAM — eager (GB) 14.52 15.00 15.17 15.15
+Latency — SDPA (s) 0.08 0.16 0.16
+Peak VRAM — SDPA (GB) 14.52 15.15 15.17
+Latency — FlashAttn (s) 0.08 0.15 0.15
+Peak VRAM — FlashAttn (GB) 14.52 15.00 15.17
+Table 7: Inference cost of different decoding strategies on LLaVA-1.5-7B. Dashes indicate that our current ASCD prototype can be integrated only in the eager path.
+Observation.
+In 7, ASCD adds at most $0.01$ s latency and $\sim$0.6 GB VRAM over the unmodified model, while VCD/ICD incur larger overhead due to image perturbations or prompt inflation.
+Thus, the selective head-level edits of ASCD preserve virtually the same inference efficiency, confirming that attention steering is a practical choice for real-time deployment.
+Appendix H Ablation Study of Hyperparameters
+Hyper-parameter robustness.
+Table 6 sweeps the four controllable hyper-parameters of ASCD—positive-steer weight $\alpha_{\text{pos}}$, contrastive weight $\alpha$, truncation threshold $\beta$, and the number of text-centric heads $\kappa_{\text{TCH}}$—on the POPE benchmark with LLaVA-1.5-7B, whose mean accuracy and F1 jointly capture hallucination-reduction and VQA quality.
+Across the entire search range each variant consistently surpasses the untuned baseline, and the performance curve is flat around the optima (bold).
+The result indicates that ASCD is insensitive to moderate hyper-parameter changes and can be deployed with minimal tuning effort.